### v1

# Multi-objective Optimization in CPU Design Space Exploration: Attention is All You Need

Abstract.
Design space exploration (DSE) enables architects to systematically evaluate various design options, guiding decisions on the most suitable configurations to meet specific objectives such as optimizing performance, power, and area. However, the growing complexity of modern CPUs has dramatically increased the number of micro-architectural parameters and expanded the overall design space, making DSE more challenging and time-consuming. Existing DSE frameworks struggle in large-scale design spaces due to inaccurate models and limited insights into parameter impact, hindering efficient identification of optimal micro-architectures within tight timeframes.
In this work, we introduce AttentionDSE [^1] The code is now available on anonymous GitHub.. Its key idea is to use the attention mechanism to establish a direct mapping of micro-architectural parameters to their contributions to predicted performance. This approach enhances both the prediction accuracy and interpretability of the performance model. Furthermore, the weights are dynamically adjusted, enabling the model to respond to design changes and effectively pinpoint the key micro-architectural parameters/components responsible for performance bottlenecks. Thus, AttentionDSE accurately, purposefully, and rapidly discovers optimal designs. Experiments on SPEC 2017 demonstrate that AttentionDSE significantly reduces exploration time by over 80% and achieves 3.9% improvement in Pareto Hypervolume compared to state-of-the-art DSE frameworks while maintaining superior prediction accuracy and efficiency with an increasing number of parameters.
1. Introduction
Design space exploration (DSE) is an essential phase in CPU design, characterized by an iterative exploration that focuses on balancing performance, power, and area (PPA) (PPA1, ; PPA2, ; PPA3, ; PPA4, ). Typically, the DSE can be formulated as a Bayesian optimization (BO) process to find multi-objective trade-offs. A prevailing DSE framework primarily comprises two components: the surrogate model and the acquisition function. Specifically, the surrogate model predicts the PPA of micro-architectural parameters, bypassing time-consuming simulations (PPA2, ; blackbox2, ; blackbox3, ; BOOMExplorer, ; MoDSE, ), and the acquisition function identifies candidate optimal alternatives for multi-objective optimization, guiding the DSE iterations (nonblackbox1, ; nonblackbox2, ; nonblackbox3, ; nonblackbox4, ; ActBoost, ; bottleneck_analysis1, ).
Concurrently, the growing complexity of modern CPUs (PowerPC_cpu, ; AMD_cpu, ; superscalar_cpu, ) results in an exponential increase in micro-architectural parameters, known as the high-dimensional design space. High-dimensional DSE imposes significant limitations on current DSE frameworks, making it difficult to identify optimal designs within a reasonable timeframe, with some even failing to find optimal solutions altogether. Specifically, it presents three key challenges.
  • $\bullet$
    First, the effectiveness of DSE frameworks is constrained by the limited accuracy and speed of surrogate models in high-dimensional design spaces. To expedite DSE, various statistical regression methods (ActBoost, ; MoDSE, ; BO1, ; BO2, ; BOOMExplorer, ; Archranker, ) have been proposed as surrogate models to predict the performance metrics of micro-architecture. However, these models suffer from an exponential increase in training and inference time (random_forest, ; GPR, ) and their accuracy decreases with the expanding of the design space (dimgpr, ; dimencurse, ).
  • $\bullet$
    Second, the effectiveness of DSE frameworks is limited by the inefficiency of acquisition functions in high-dimensional design spaces. Acquisition functions are crucial in the DSE framework, as they identify the candidate optimal design for exploration. However, most acquisition functions depend heavily on expert knowledge to identify the most potential sample in the design space (ActBoost, ; MoDSE, ; BOOMExplorer, ), or on complex feature engineering (ArchExplorer, ) which requires a time-consuming process to identify the bottleneck of the current micro-architecture. Moreover, some of them even depend on exhaustive search (MoDSE, ; BOOMExplorer, ; ActBoost, ), which makes the acquisition functions impractical in high-dimensional design spaces.
  • $\bullet$
    Third, recent DSE frameworks overlook the impact of individual micro-architectural parameters on final performance metrics, missing valuable insights into their contributions. This issue primarily arises from the surrogate model’s inability to accurately capture the contribution of each parameter to overall performance (MoDSE, ; BOOMExplorer, ; ActBoost, ; PPA1, ; challenge3_3, ). Additionally, the separation of the surrogate model and acquisition function as independent entities in previous efforts further exacerbates this issue, as it impedes the seamless information exchange between them (ArchExplorer, ; MoDSE, ; BOOMExplorer, ; ActBoost, ).
In this work, we introduce AttentionDSE, a novel DSE framework that leverages the attention mechanism to accurately, purposefully, and rapidly identify optimal CPU designs in the high-dimensional design space. The key idea of AttentionDSE lies in its use of attention-based models to establish a direct mapping between micro-architectural parameters and performance metrics. This approach improves both the accuracy and interpretability of the performance model. Additionally, the attention weights are dynamically adjusted throughout the exploration process, allowing the model to adapt efficiently to design changes. This dynamic adjustment enables more effective identification of key micro-architectural parameters responsible for performance bottlenecks. As a result, AttentionDSE not only accelerates the design space exploration process but also ensures more precise and purposeful optimization, enabling rapid discovery of optimal micro-architectural designs.
Overall, our contributions are as follows:
  • $\bullet$
    We transform the DSE into a sequence prediction problem, introducing AttentionDSE that synergizes the surrogate model and acquisition function and leverages the attention mechanism for high-dimensional DSE on CPU.
  • $\bullet$
    We introduce a Perception-Driven Attention (PDA) mechanism, which utilizes micro-architectural insights to focus on relevant parameters, minimizing computational overhead of irrelevant parameters to reduce the training and inference time of attention-based performance model.
  • $\bullet$
    We propose an Attention-aware Bottleneck Analysis (ABA) algorithm, which uses attention weights to purposefully and quickly pinpoint bottlenecks, guiding the exploration.
  • $\bullet$
    Experimental results show that AttentionDSE reduces exploration time by over 80% and improves Pareto Hypervolume (PHV) by 3.9% compared to state-of-the-art (SOTA) DSE frameworks. Additionally, as the number of micro-architectural parameters increases, AttentionDSE keeps stable prediction accuracy and avoids significant increases in training and inference time.
2. Background
In this section, we introduce the relevant concepts.
2.1. Attention Mechanism
The attention mechanism has become one of the most crucial concepts in deep learning. Inspired by human biological systems, which focus on distinctive and relevant details when processing large amounts of information, attention mechanisms allow models to prioritize important data instead of treating every input equally. This selective focus enhances the efficiency and effectiveness of handling complex data.
A typical attention layer in a neural network dynamically identifies the most relevant parts of the input. The attention score is defined as:
(1) $\operatorname{Attention}(Q,K,V)=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V.$
It operates using three key vectors: query (Q), key (K), and value (V). The attention mechanism calculates a score by comparing the query and key vectors, which are normalized via a softmax function to produce attention weights. These weights then compute a weighted sum of the value vectors, emphasizing the most critical information. This process is especially effective in self-attention, where each input element attends to all others, and in multi-head attention, which captures different aspects of the input through multiple attention heads. Attention layers are foundational in models like transformers (transformer, ), enabling flexible and efficient learning of dependencies within sequences.
2.2. Design Space Exploration in CPU Design
DSE is a critical technique in the optimization and design process, primarily used to identify and analyze various design options within CPU design (PPA1, ; PPA2, ; PPA3, ; PPA4, ). Its goal is to systematically explore all possible design alternatives to find the optimal or near-optimal solutions that meet specific design requirements and constraints (ArchExplorer, ; MoDSE, ; BOOMExplorer, ; ActBoost, ; wangduo1, ; wangduo2, ).
To expedite the process, prevailing DSE frameworks adhere to the fundamental principles of BO (MoDSE, ; BOOMExplorer, ; ActBoost, ; Archranker, ). BO leverages a surrogate model that serves as a predictive tool, approximating the underlying objective function (GPR, ; random_forest, ; ensemble_learning, ). Additionally, an acquisition function drives the optimization process in DSE frameworks, efficiently navigating the DSE by identifying promising design points (PPA2, ), which is a set of micro-architectural parameters, for evaluation (ArchExplorer, ). The above two facilitate effective exploration of the design space within a reasonable time (BO1, ; BO2, ).
Figure 1. The workflow of prevailing DSE framework. ❶~❹ form the exploration loop.
Fig. 1 illustrates the core workflow of the BO-based DSE framework. Prior to the exploration stage, the framework first trains the surrogate model using sampling methods, such as Plackett-Burman ranking(PBrank, ; MoDSE, ), to generate the training set (①, ②, ③). Once the surrogate models are trained, the exploration stage begins. The acquisition function selects a design point from the design space (❶) and sends it to the surrogate models to evaluate the PPA (❷). The framework then iteratively updates the Pareto-optimal set and repeats this process (❸, ❹).
2.3. Pareto Optimal Set
A key concept in CPU DSE is the Pareto optimal set. A solution is considered as Pareto optimal if there is no other solution that improves some objective without worsening at least one other objective. To formalize the DSE as a multi-objective optimization task, the micro-architectural parameters are serialized as a feature vector $x$ referred to as a design point. The collection of all design points forms the design space $D$. Performance Metrics used to evaluate a design point are denoted as $y=f(x)$. In an n-objective minimization problem, a design point $x$ is defined to be dominated by $x^{*}$ if
(2) $\forall i\in[1,n],f_{i}\left(x^{*}\right)\leq f_{i}(x);\exists j\in[1,n],f_{j}\left(x^{*}\right)<f_{j}(x).$
This dominance is denoted as the partial order $x^{*}\prec x$ and $y^{*}\prec y$. Conversely, if $x^{*}$ does not dominate $x$, it is denoted as $x^{*}\nprec x$ and $y^{*}\nprec y$. Among all design points, the set of design points that are not dominated by any other design points is called the Pareto optimal set $\Omega$, formulated as:
(3) $\Omega=\left\{x\mid x^{*}\nprec x,\forall x^{*}\in D\right\}.$
The PHV is determined by the Pareto optimal set and a reference point, as shown in Fig.2(a). It is represented by the region of blue points, enclosed by light gray lines, with the reference point labeled as $Z_{\text{ref}}$. A larger PHV indicates a better compromise achieved by the solution set across multiple optimization objectives. In Fig.2(b), the new design points $x$ and $y$ demonstrate superior performance or power compared to the original design points, with both expanding the PHV, illustrated by the red and purple regions, respectively.
Figure 2. The Expansion of PHV: (a) original PHV; (b) design points excelling in performance or power extend the PHV.
3. Related Works
For the early work, Mariani et al. (relative_work1, ) use a Kriging model (Kriging, ) with exponential correlation as a surrogate to predict the Pareto nondominated rank distribution. They then iteratively optimize using the Expected Improvement (EI) of non-dominated levels as the acquisition function. CASH (relative_work2, ) compares multiple machine learning models and finds that BO achieves superior exploration results, by using a random forest as the surrogate model and EI as the acquisition function. Wang et al. (relative_work3, ) employ an adaptive component selection and smoothing operator as the surrogate model and the expected Hypervolume Improvement (HVI) of the Pareto frontier as the acquisition function.
For the blackbox DSE framework, ActBoost (ActBoost, ) integrates statistical methods and machine learning to prune the design space and determine the training set. For the surrogate model, ActBoost uses two AdaBoost.RT (AdaBoost.RT, ) models with active learning. For the acquisition function, it calculates the prediction uncertainty of each unsampled design point and selects one with the highest uncertainty. By combining statistical sampling and the boosting technique, ActBoost enables efficient and accurate DSE tasks. BOOMExplorer (BOOMExplorer, ) employs Gaussian Process Regression (GPR) (GPR, ) as the surrogate model and leverages EI (EI, ) as the acquisition function, searching for the next design point by analyzing data distribution characteristics. By embedding prior knowledge, BOOMExplorer achieves superior compared to analytical methods. MoDSE (MoDSE, ) leverages ensemble learning for more accurate prediction. It combines AdaBoost (AdaBoost, ) and Gradient Boosted Regression Trees (GBRT) (GBRT, ) as the surrogate model and uses HVI with a full-space search for the acquisition function. MoDSE introduces a uniformity-aware selection algorithm for efficient multi-objective DSE and a Pareto-rank-based sample weight generation algorithm to train the surrogate model, enhancing the DSE outcomes.
For the DSE frameworks based on bottleneck analysis, ArchExplorer (ArchExplorer, ) adopts the dynamic event-dependence graph (DEG) (DEG, ) for critical path analysis and identifies bottlenecks in the current micro-architecture to guide the design of micro-architecture. It couples this approach with a software simulator (Gem5) (gem5, ) as the surrogate model. ArchExplorer focuses on refining the acquisition function to reduce the need for domain knowledge in mechanistic models. Explainable-DSE (Explainable_DSE, ) is the SOTA framework for DSE tasks in deep neural network (DNN) accelerators. By leveraging the structure and characteristics of neural networks, Explainable-DSE constructs a bottleneck model to identify design bottlenecks. It enables targeted optimizations to mitigate issues such as high latency and power consumption, while also providing explanations for the design adjustments made and their impact on performance. This framework is primarily used for the joint DSE of DNN accelerators and DNN algorithms.
4. Motivation
In this section, we will explain the motivations behind leveraging the attention mechanism in the DSE framework.
4.1. Challenges in Current DSE Studies
Modern CPU design is becoming increasingly complex, with a growing number of micro-architectural parameters. The exploration in such a significantly large design space is known as the high-dimensional DSE. Current DSE frameworks face three primary challenges in high-dimensional tasks.
First, DSE frameworks struggle with effectiveness due to the limited accuracy and efficiency of surrogate models in high-dimensional design spaces. The surrogate models in prevailing DSE frameworks (ActBoost, ; MoDSE, ; BOOMExplorer, ; Archranker, ), which typically use statistical regression methods, are simple while vulnerable. Specifically, prediction accuracy can decline as dimensionality increases. For instance, when the number of parameters exceeds 75, BOOMExplorer struggles to accurately predict and associate micro-architectural parameters with performance metrics.
This issue arises from the sparser feature space in high-dimensional condition (dimencurse, ), indicating the need for more training data to accurately capture the data distribution. However, the training and inference time increases greatly as the size of the training set grows. For instance, with a training set of size $n$, the complexity of training the inverse of the covariance matrix typically reaches $O(n^{3})$ and the inference through the inverse of the covariance matrix is $O(n^{2})$. This considerable computational cost renders high-dimensional design spaces impractical.
Second, the performance of DSE frameworks is hindered by the inefficiency of acquisition functions when dealing with high-dimensional design spaces. Acquisition functions are crucial for finding optimal design configurations, but they face challenges in high-dimensional design spaces. Current methods often rely on expert knowledge (ActBoost, ; MoDSE, ; BOOMExplorer, ) or complex feature engineering (ArchExplorer, ) to identify the bottleneck of the current micro-architecture, which can introduce biases, require significant time, and lack scalability. For instance, ArchExplorer only analyses the first hundred thousand of the instructions and achieves the optimal architecture with the number of parameters 21 over 15 days (ArchExplorer, ). Moreover, some methods rely on large-scale or even exhaustive search (MoDSE, ; ActBoost, ; BOOMExplorer, ). In high-dimensional spaces, this exhaustive search becomes computationally prohibitive due to the enlarged design space. For example, MoDSE can only perform with the number of parameters 10 (totaling 36,864 design points) (MoDSE, ). When the number of parameters increases to 30, it may take $1.9\times 10^{29}$ years. The high computational cost of such searches slows down the DSE process, delaying the identification of high-performance designs. This inefficiency slows down the DSE process, highlighting the need for more advanced and efficient acquisition strategies to handle complex, high-dimensional problems.
Third, conventional DSE frameworks fail to recognize the impact of each parameter on performance metrics. This limitation can cause these frameworks to overlook critical insights into how individual parameters impact the performance metrics, making it difficult to quickly and accurately identify bottlenecks in the current design. This phenomenon can be attributed to two primary reasons. Firstly, existing surrogate models cannot unveil the contributions of each parameter to the performance metrics because they rely on mathematical analysis (ActBoost, ; Archranker, ) or simplistic black-box models (BOOMExplorer, ; MoDSE, ) for prediction. Secondly, prevailing DSE frameworks typically adhere to a two-stage DSE scheme, treating the surrogate model and acquisition function as independent entities. This separation inhibits the exchange of information, preventing a comprehensive understanding of parameter efficacy (BOOMExplorer, ; MoDSE, ; ActBoost, ; ArchExplorer, ). Moreover, the lack of integration between these two components hinders the ability of the acquisition function to adaptively refine the exploration results based on feedback from the surrogate model, resulting in suboptimal performance and less effective DSE.
4.2. Opportunity of Attention in DSE Tasks
DSE is a crucial process in CPU design, often consisting of two important tasks: prediction and bottleneck analysis. Prediction helps estimate the performance of various design configurations, while bottleneck analysis identifies critical constraints that limit system efficiency. Both tasks are essential to guide CPU design effectively.
Recent advances in attention mechanisms offer significant advantages in addressing these challenges, proving to be highly effective for both performance prediction and bottleneck analysis. First, attention has been demonstrated to excel in regression tasks (attention_regression1, ; attention_regression2, ), accurately capturing complex relationships between inputs and outputs. Second, attention mechanisms inherently highlight the importance of different parameters by assigning varying weights, thus enabling a more fine-grained analysis of key design factors. These strengths align directly with the core needs of DSE, where both accurate performance prediction and the identification of critical design parameters are essential for driving effective design exploration.
5. Design of AttentionDSE
In this section, we introduce AttentionDSE, the first DSE framework that adopts the attention mechanism, suitable for high-dimensional DSE.
5.1. Overview of AttentionDSE
Figure 3 provides an overview of AttentionDSE. It comprises three primary components: 1) Predictors, which provide rapid and accurate predictions of performance metrics (e.g., IPC) for given micro-architectures across benchmarks, thereby obviating the need for time-intensive simulations; 2) Micro-architecture Serialization, designed to reduce the computational burden associated with predictions, thereby expediting the performance evaluation process; 3) Bottleneck Analysis and Iterator, which identify the micro-architectural parameters that introduce performance bottlenecks within micro-architecture modules, facilitating the exploration of promising micro-architectural candidates by selectively modifying these parameters. The process of the AttentionDSE consists of a predictor training stage (①, ②, ③, ④) and an architecture exploration stage (❶, ❷, ❸, ❹, ❺).
Figure 3. Overview of the AttentionDSE.
Figure 4. The workflow of AttentionDSE framework in the exploration stage.
The training stage is a one-time process, and the sequence of operations is represented by hollow numbers. This stage focuses on training predictors for performance metrics, ensuring that the models accurately capture the relationships between micro-architectural parameters and performance metrics. Unlike other approaches that rely on specialized sampling methods, AttentionDSE adopts a random sampling approach and then converts the design point to the sequence input, generating the training set (①, ②, ③). Next, the deep learning model, integrated with the PDA, functioning as the predictor, is trained to predict performance metrics (④). Benefiting from probabilistic attribute correlation in attention calculation, it’s more convenient to capture the intricate relationship between parameters, which informs exploration hints for the exploration stage.
Once the predictors are trained, AttentionDSE is ready for the exploration stage, which is iterative and aims for efficient multi-objective optimization. The exploration stage begins with a random sampling of design points (❶), followed by the exploration of the Pareto optimal set of these points through the trained predictors (❷). Each design point in the Pareto optimal set is analyzed to determine the contribution of each parameter (❸). The Bottleneck Analysis Iterator then selects the next design point based on the ABA algorithm (❹, ❺). This iterative process continues until the preset performance metrics are met, ensuring efficient and thorough exploration of the design space. Fig. 4 presents a detailed workflow of the exploration stage. All these techniques will be discussed in detail in the following sections.
5.2. Attention-based Prediction Model
A fast and accurate performance predictor can replace traditional simulation methods, serving as the cornerstone of an efficient DSE framework. This approach significantly accelerates design iterations and reduces the overall design cycle. As discussed in Section 4, the attention-based model is a promising option for developing such a predictor. Neural networks possess strong nonlinear fitting capabilities, making them well-suited for learning from high-dimensional data. Additionally, they benefit from robust hardware acceleration and a well-developed GPU-based framework, collectively enhancing training and inference speed.
In this subsection, we propose an attention-based prediction model as the single-objective predictor. We aim to develop a more accurate and scalable predictor that emphasizes efficiency in high-dimensional DSE.
Input: $D$: Design space; $n$: The size of the training set; $k$: Embedding length; $Depth$: The depth of the model; $Epoch$: The total epochs for training.
Output: $M$: The trained predictor.
1
[…]
3 Instantiate the prediction token $P_{t}$ with embedding length $k$;
4 Initialize the weights of the Predictor;
5 Simulate $T$ to obtain $S_{ground\_truth}$ of IPC, Power, and Area;
6 for $epoch\leftarrow 0$ to $Epoch$ do
7       $S\leftarrow Serialize(T)$;
8       $S_{embed}\leftarrow Embedding(S)$;
9       $S_{embed}\leftarrow Concat(S_{embed},P_{t})$;
10       for $depth\leftarrow 0$ to $Depth$ do
11             $S_{attn}\leftarrow SelfAttention(S_{embed})$;
12             $S_{mlp}\leftarrow MLP(S_{attn})$;
13
14       end for
15      $S_{attn},A_{score}\leftarrow SelfAttention(S_{mlp})$;
16       Extract processed prediction token $S_{attn}\leftarrow S_{attn}[0,:]$;
17       $S_{pred}\leftarrow FullyConnectedLayer(S_{attn})$;
18       $BackwardPropagation(S_{pred},S_{ground\_truth})$;
19
20 end for
21return Predictor;
Algorithm 1 The Training Procedure of Predictors
The training procedure of the attention-based prediction model is demonstrated in Algorithm. 1. The prediction model consists of two key components: embedding preprocessing and the attention-based model. For embedding preprocessing, the procedure involves three steps: (a) Embedding, which transforms the discrete and heterogeneous alternatives of each micro-architectural parameter into a continuous numerical representation, simplifying optimization and enhancing the representation of relationships within the micro-architecture; (b) Position Embedding (PE), which assigns positional encoding to each item in the parameter sequence, enabling the model to learn parameter attributes more effectively and constrain the embedding space of each parameter; and (c) Prediction Token (PT), which gathers information from parameter embeddings using the self-attention mechanism and fuses it through multilayer perceptron (MLP), allowing the model to perceive the entire parameter sequence effectively. As for the attention-based model, to ensure the generalizability and versatility of the model, the predictor includes only vanilla self-attention, MLP, and normalization operations, which are the minimal necessary components of the transformer model (transformer, ). This streamlined design aims to maintain the attention mechanism’s inherent advantages. By focusing on these core components, we can leverage the self-attention mechanism’s ability to capture information within the input design point, allowing for the effective handling of complicated relationships between the design point and performance metrics in high-dimensional DSE tasks. The inclusion of MLP layers provides the model with the capacity to learn complex patterns and representations, while normalization operations ensure stable and efficient training by mitigating issues related to internal covariate shifts.
5.3. Perception-driven Attention Mechanism
Although the prediction time is significantly reduced compared to simulation, further reductions are still necessary, especially when handling the large-scale design spaces of CPUs. Fortunately, from a micro-architectural design perspective, not all micro-architectural parameters are strongly correlated with one another, presenting an opportunity to reduce the computational burden of the attention mechanism significantly. For instance, the fetch buffer is closely related to the fetch width and fetch queue but has only a weak relationship with the writeback width. Therefore, in this section, we introduce the PDA mechanism as an alternative to the vanilla self-attention mechanism to alleviate the computational burden. The PDA mechanism includes two aspects: the micro-architecture serialization method and the sliding window attention mechanism.
In previous work (GPR, ; MoDSE, ; ensemble_learning, ; BOOMExplorer, ), the order of parameters is not a concern. Because GPR (GPR, ; BOOMExplorer, ) relies on the distance and similarity between input data rather than their arrangement order in statistical regression methods. Similarly, in ensemble learning (ensemble_learning, ; MoDSE, ), the training process of multiple weak learners introduces randomness, making the order of parameters irrelevant. To address this challenge in the attention-based predictor, we introduce a serialization approach named Perception-driven Serialization (PDS), aimed at preserving architectural information and laying the groundwork for the sliding window attention technique. This technique employs a graph-based method to delineate relationships among all parameters within each pipeline stage.
Figure 5. The Perception-driven Serialization approach. (a) A Sample OoO CPU core; (b) Rename Stage; (c) Stage abstraction; (d) The steps of Perception-driven Serialization approach.
An example is illustrated in Fig. 5(a). We chose the Rename stage, shown in Fig. 5(b), as an example, where there are four parameters to be serialized including the free list (F), the rename map (R), the scoreboard (S), and the history buffer (H). The PDS first transfers the micro-architecture into a perceptual graph, which is an undirected graph based on the dataflow among these components, as depicted in Fig. 5(c). In a perceptual graph, vertices represent the parameters within each pipeline stage, while edges denote the datapaths between these components. The edges are categorized into two types: red edges, which connect to other pipeline stages and are labeled as external edges, and black edges, which connect to components within the current pipeline stage and are labeled as internal edges. To quantify the perception of each component, we leverage the perception degree $D$ of each parameter in the perceptual graph. The perception degree is calculated as follows:
(4) $D(\text{parameter})=\sum_{e\in E_{\text{parameter}}^{\text{internal}}}e-\sum_{e\in E_{\text{parameter}}^{\text{external}}}e$
Here, the internal edges represent the intra-stage perception field, while the external edges represent the inter-stage perception field. After calculating the perception degrees, the sequence of the micro-architecture is determined using the insertion sort algorithm. Parameters with the larger perception degree are inserted into the middle of the sequence as shown in Fig. 5(d). After all the pipeline stages have been serialized, the final vector of the micro-architecture is combined according to the order of the pipeline stages.
Figure 6. The attention weight calculation. (a) Vanilla self-attention mechanism; (b) PDA mechanism.
After the micro-architecture has been serialized, the vanilla self-attention mechanism can be replaced by the sliding window attention mechanism due to the parameter clusters. Fig. 6 shows an example of attention weight calculation compared with vanilla self-attention and PDA. For vanilla self-attention, every parameter must calculate the attention weight with every other parameter, as shown in Fig. 6(a). To optimize this process, we set the window size to the maximum perception degree among all parameters. In this example, the window size is set to five, as illustrated in Fig. 6(b).
In general, with the help of the PDA, each parameter calculates the attention weight only with its relative parameters. In this way, for a parameter vector of length $n$, the computational complexity is reduced from $O(n^{2})$ to $O(n)$. Thus, the PDA approach alleviates the limitations posed by number of parameters, enhances scalability, and reduces the training and inference time of the prediction model.
5.4. Attention-aware Bottleneck Analysis
Bottleneck analysis is a critical step in CPU design, enabling architects to identify the key factors that impact system performance. In this subsection, we propose the ABA algorithm, which equips AttentionDSE with the capability to analyze attention weight heatmaps generated by an attention-based predictor. The ABA algorithm, detailed in Algorithm 2, reveals the intricate relationships between micro-architectural parameters and performance metrics, thus guiding the exploration of design space. This approach ensures that the most influential design parameters are prioritized during optimization, leading to more efficient and effective DSE.
Input: $D$: The design space; $X$, $X_{n}$: Initial design point set and its size; $I_{max}$: The max iteration of exploration;
Output: $\Omega$: The Pareto optimal set
1
[…]
3 $\Omega\leftarrow GetParetoOptimal(Predictor(X))$;
4 for $i$leftarrow 0 to $I_{max}$ do
5       Set search queue $Q\leftarrow\varnothing$;
6       for each design point x in $\Omega$ do
7             $x^{\prime}\leftarrow BottleneckAnalysis(x)$;
8             $PPA\leftarrow Predictor(x^{\prime})$;
9             if $x^{\prime}~{}has~{}expended~{}the~{}\Omega$ then
10                   $Q.push(x^{\prime})$;
11
12             end if
13
14       end for
15      $\Omega\leftarrow GetParetoOptimal(Predictor(Q))$;
16
17 end for
18return $\Omega$;
Algorithm 2 Attention-aware Bottleneck Analysis
In the attention weight heatmap, each row represents how the current parameter is influenced by other parameters, reflecting the correlations between them. For a particular parameter, the sum of all weights in its corresponding column indicates the overall influence of this parameter on others. Therefore, if the total sum of attention weights in the column corresponding to a certain parameter is minimal for IPC or maximal for Power or Area, we infer that this parameter has a significant impact on the current performance metrics. ABA algorithm then updates the Pareto optimal set by updating the design point based on this bottleneck analysis, ensuring continuous improvement and exploration of the design space.
Figure 7. The example of ABA algorithm updating the micro-architectural parameters.
To clarify the execution process of the ABA algorithm, Fig. 7 illustrates a toy example that demonstrates how the ABA algorithm updates micro-architectural parameters step by step to enhance performance (IPC). Initially, AttentionDSE randomly samples a design point and sends the parameters to the predictor. Based on the attention weight heatmap from the self-attention layer, the ABA algorithm sums all the columns and selects the parameter with the minimal sum as the current performance bottleneck, which in this example is the integer ALUs. The ABA algorithm then increases the number of ALUs from six to eight, improving performance. In the subsequent iteration, the ABA algorithm identifies the instruction queue size as the current bottleneck using the same process. The ABA algorithm increases the queue size from 64 to 80. By strategically adjusting the instruction queue size, the pipeline is provided with a sufficient flow of instructions, further optimizing performance.
6. Experimental Setup
In this section, we introduce the experimental setup used to assess the performance of AttentionDSE.
Simulation. We utilize GEM5 (gem5, ) as the timing-accurate simulator and McPAT (mcpat, ) as the power and area modeling tool. We extend GEM5 using Python to generate different CPU cores with varying configurations. To bolster the credibility of our experiments, we employ SPEC CPU2017 (SPEC2017, ) as the evaluation benchmark. For identifying resource utilization status, we utilize Simpoints (simpoints, ) for each workload evaluation. Each workload is divided into at most 30 clusters, with each cluster containing ten million instructions.
Baselines. Our baselines include the ActBoost (ActBoost, ), the BOOMExplorer (BOOMExplorer, ), and the MoDSE (MoDSE, ). Additionally, we implement ArchExplorer (ArchExplorer, ) for comparison with our ABA algorithm. The baselines represent recent SOTA work. All experiments are conducted on the Out-of-Order (OoO) CPU implemented by Gem5. The design space of the OoO processor is listed in Table 1. The values in the third column are formatted as “start number: end number: stride”. A two-level cache hierarchy with an 8192MB DRAM main memory configuration is employed. The size of the design space exceeds $6.89\times 10^{35}$, making exhaustive search impractical.
Table 1. Microarchitecture design space specification.
Parameters Description
Candidate
Value
Core Frequency
the frequency of
CPU core in GHz
1/1.5/2/
2.5/3
Pipeline Width
fetch/decode/rename/
dispatch/issue/writeback/
commit width
1:12:1
Fetch Buffer
fetch buffer size
in bytes
16/32/64
Fetch Queue
fetch queue size
in $\mu$-ops
8:48:4
Branch Predictor predictor type
BiModeBP/
TournamentBP
Choice Predictor choice predictor size 2048/4096/8192
Global Predictor global predictor size 2048/4096/8192
RAS Size return address stack size 16:40:2
BTB Size branch target buffer size 1024/2048/4096
ROB Size reorder buffer entries 32:256:16
Int RF Number
number of physic
integer registers
64:256:8
Fp RF Number
number of physical
floating-point registers
64:256:8
Inst Queue
number of instruction
queue entries
16:80:8
Load Queue
number of
load queue entries
20:48:4
Store Queue
number of
store queue entries
20:48:4
IntALU number of integer ALUs 3:8:1
IntMultDiv
number of integer
multipliers and dividers
1:4:1
FpALU
number of
floating-point ALUs
1:4:1
FpMultDiv
number of floating-point
multipliers and dividers
1:4:1
Cacheline cacheline size 32/64
L1 ICache Size size of ICache in KB 16/32/64
L1 ICache Assoc.
associative sets of
ICache
2/4
L1 DCache Size size of DCache in KB 16/32/64
L1 DCache Assoc.
associative sets of
DCache
2/4
L2 Cache Size size of L2 Cache in KB 128/256
L2 Cache Assoc.
associative sets of
L2 Cache
2/4
Total size $6.89\times 10^{35}$
Evaluation Metrics. A DSE framework is evaluated mainly by the accuracy of the predictor and the quality of the final Pareto optimal set.
The accuracy is mainly measured by Mean Absolute Percentage Error (MAPE), which calculates the average absolute percentage difference between real and predicted values. The formula for MAPE is:
[…]
where $n$ is the number of predicted design points, $y_{i}$ is the prediction objective value, i.e., IPC, power, or area in corresponding models, and $y_{i}^{real}$ is the real value; MAPE expresses the prediction error as a percentage of the actual values. A lower MAPE indicates better accuracy.
The PHV metric assesses the quality of the Pareto optimal set by quantifying the volume of the objective space dominated by non-dominated solutions. The PHV can be formulated as:
(6) $\mathrm{PHV}(\Omega)=\int_{\mathrm{X}}\mathbb{1}[\boldsymbol{x}\prec\boldsymbol{x_{ref}}][1-\prod_{\boldsymbol{x}^{*}\in\Omega}\mathbb{1}[\boldsymbol{x}^{*}\nprec\boldsymbol{x}]]\mathrm{d}\boldsymbol{x}.$
Where $\mathbb{1}(\textperiodcentered)$ is the indicator function, which outputs one if its argument is true and zero otherwise, specifically, if $x$ is the newly discovered design that is not dominated by other design points, then the PHV increases. An example is illustrated in Fig.2.
7. Experiment Results
7.1. Overall Analysis
We evaluate the AttentionDSE framework using SPEC CPU 2017 workloads.
7.1.1. Comparison on Pareto Optimal Set Exploration
Figure 8. Comparison of PHV from IPC-Power optimization with SOTA DSE frameworks.
Figure 9. Comparisons between the Pareto designs in (a) $Perf^{2}/(Power\times Area)$, (b) performance, (c) area, and (d) power.
Fig. 8 provides a more detailed comparison of the Pareto optimal set exploration through different DSE frameworks. Compared to SOTA DSE frameworks, AttentionDSE achieves an average improvement in PHV of 8.3%, 4.1%, and 3.9% across all workloads over ActBoost, BOOMExplorer, and MoDSE, respectively. ArchExplorer employs a similar concept to AttentionDSE, analyzing program bottlenecks and addressing them accordingly. It uses DEG to identify micro-architecture bottlenecks, a process that can take several days. AttentionDSE outperforms ArchExplorer with a 1.3% improvement in PHV. This superiority is due to two main factors: First, AttentionDSE uses an attention mechanism for design point updates and bottleneck identification, accurately pinpointing issues in the architecture. Second, ArchExplorer’s bottleneck analysis is limited by complexity constraints, as it only examines the first hundred thousand lines of code. This partial analysis fails to provide a comprehensive understanding of the program, leading to reduced performance. Note that MoDSE’s exploration using the HVI approach, which searches the entire set to find the next design point, is feasible only in small design spaces. In this experiment, we use a random sampling method for design point search in MoDSE.
Table 2. Detailed information of Pareto optimal set search.
Methods
[…]
PHV
Up to 99%
PHV Iterations
Time
(Minutes)&(Ratio)
BOOMExplorer 36.94 64 47 (9.4$\times$)
MoDSE 37.04 87 16 (3.2$\times$)
ArchExplorer 38.57 \ 828 (165.6$\times$)
AttentionDSE 39.41 15 5 (1$\times$)
Note: The experimental results are obtained from IPC-Power optimization on the 600.perlbench_s dataset.
Table 2 shows the exploration results on the 600.perlbench_s dataset across the SOTA DSE frameworks. The “Time” represents the total exploration time of the DSE task. All DSE frameworks, except ArchExplorer, perform a hundred iterations for the Pareto optimal set search. The “Up to 99% PHV Iteration” metric indicates the convergence time of the search, where lower values are better. Experimental results demonstrate that AttentionDSE is 166 $\times$ faster than ArchExplorer, owing to ArchExplorer’s time-consuming instruction analysis method. Compared to regression-based models, AttentionDSE completes the exploration roughly 10$\times$ faster than BOOMExplorer. Even with MoDSE’s use of random sampling, AttentionDSE achieves exploration speeds approximately 3$\times$ faster. Additionally, AttentionDSE converges in fewer iterations, reducing iterations by 75% compared to BOOMExplorer and 83% compared to MoDSE, highlighting the efficiency of the ABA algorithm.
Overall, the results demonstrate that AttentionDSE provides significant improvements over existing work, validating the effectiveness of its novel methodologies in DSE.
Figure 10. Comparison of MAPE with SOTA DSE frameworks. (a) MAPE of IPC prediction; (b) MAPE of power prediction; (c) MAPE of area prediction; (d) MAPE and normalized inference time with the increased number of parameters on the 600.perlbench_s dataset; (e) The training time changes with the increase in training set size on the 600.perlbench_s dataset.
7.1.2. Comparison on Best Balanced Designs
To study how high-performance designs balance PPA effectively, we select the designs with the highest performance for comparison. The metric we used for this evaluation is $Perf^{2}/(Power\times Area)$, which assesses how well each design balances performance with resource constraints. Fig. 9 gives the results for SPEC CPU 2017, comparing frameworks such as ActBoost, BOOMExplorer, MoDSE, ArchExplorer, and AttentionDSE, with the last bar representing the geometric mean.
Across all metrics, AttentionDSE achieves, on average, the highest performance while maintaining the lowest energy consumption and area overhead, demonstrating the superiority of our methodology. Specifically, as shown in Fig. 9(a), AttentionDSE outperforms the other methods by an average of 13.8%, 6.9%, 7.8%, and 5.6%. Although AttentionDSE does not achieve the highest performance for every workload, as shown in Fig. 9(b), it consistently maintains multi-objective optimality, as shown in Fig. 9(c) and Fig. 9(d), leading to overall optimality across the design space.
7.1.3. Comparison on Single-objective Predictor
An accurate predictor is fundamental to an effective DSE framework. Fig. 10 illustrates the comparison of the prediction models with the SOTA DSE frameworks. Fig. 10(a), (b), and (c) present the MAPE of AttentionDSE compared with three DSE frameworks. For example, AttentionDSE outperforms all workloads, achieving 37.1%, 33.4%, and 31.6% on average for IPC, Power, and Area prediction on MAPE, respectively, compared to the SOTA DSE framework (MoDSE). These results highlight the superiority of AttentionDSE over statistical regression methods. We also conduct experiments using an MLP regression model, as shown in Table 3, to demonstrate the superiority of the attention mechanism.
Table 3. Comparison of AttentionDSE with the MLP model.
[…]
IPC Power Area IPC Power Area
MAPE 0.0722 0.0399 0.0274 0.034 0.028 0.017
$R^{2}$ 0.8755 0.9689 0.9868 0.97 0.98 0.99
MSE 0.0042 0.0009 0.2510 0.00022 0.0017 0.12
Note: The experimental results are obtained on the 600.perlbench_s dataset.
To better demonstrate support for high-dimensional DSE, we carry on additional experiments on the prediction models using the 600.perlbench_s dataset with an increasing number of parameters and varying training set sizes, as shown in Fig. 10(d) and (e). In Fig. 10(d), the bar chart represents the comparison of MAPE, while the line chart indicates the inference time. BOOMExplorer, which uses GPR as the prediction model, struggles to capture effective models in high-dimensional design spaces, resulting in poor generalization performance. When the number of parameters increases to 75, BOOMExplorer fails to capture the relationship between parameters and performance metrics, resulting in a MAPE of 1. On the other hand, although MoDSE shows an acceptable range of MAPE changes, its inference time increases sharply with the number of parameters, making it impractical for high-dimensional DSE. Fig. 10(e) illustrates the training time of each prediction model as it changes with the increase in training set size. High-dimensional design spaces need more training data to capture the relationship between micro-architecture and performance metrics. However, prevailing DSE frameworks face a significant and unacceptable increase in training time as the amount of training data grows. For AttentionDSE, the training time remains stable due to several aspects: first, the predictor in AttentionDSE consists of the minimal necessary components for an attention-based model, ensuring lower training time through its lightweight design; second, AttentionDSE adopts the PDA mechanism, further decreasing computational complexity.
7.2. Optimization Effect Analysis
We further conduct an ablation experiment to demonstrate the effectiveness of the optimizations. We select 600.perlbench_s as the representative dataset, and the results are consistent across other datasets.
7.2.1. Efficiency of Perception-driven Attention Mechanism
To demonstrate the benefits of the PDA mechanism, we conduct experiments to evaluate the training time and memory footprint with varying numbers of parameters, highlighting the PDA mechanism’s superiority in high-dimensional DSE tasks. We set the model depth to ten and observe the impact on both training time and memory usage.
Figure 11. The efficiency of PDA mechanism.
Fig. 11 shows the experiment results, the line chart illustrates the normalized training time, while the bar chart depicts the memory footprint. The training time for models without the PDA mechanism increases exponentially with the size of the design space. For example, when the design space expands sixfold, the training time increases by 30$\times$. In contrast, AttentionDSE with the PDA mechanism exhibits relatively stable memory usage as the number of parameters grows. This stability extends to large design spaces because the PDA mechanism only calculates within clusters that contain the most relevant parameters. Regarding memory footprint, in smaller design spaces, memory usage is primarily influenced by fixed data occupation during training. However, attention weights and other learnable parameters in larger design spaces gradually dominate memory usage.
Figure 12. The impact of the PDA mechanism.
Fig. 12 shows the difference in MAPE introduced by the PDA mechanism. Due to the PDA mechanism clustering the relevant parameters, captured by the sliding window attention technique, the predictive accuracy does not decrease and may even slightly increase. This improvement is primarily due to the enhanced understanding of local information within the pipeline stage provided by PDA. By focusing on parameter interactions within a defined window, PDA captures intricate relationships between parameters, resulting in more accurate predictions.
7.2.2. Analysis on Attention-aware Bottleneck Analysis Algorithm
The superiority of AttentionDSE derives from both its high-accuracy prediction model and the ABA algorithm. To more effectively illustrate the impact of the ABA algorithm and avoid interference from the inaccurate prediction model, we conduct experiments on AttentionDSE by substituting the ABA algorithm with other design point search methods. Due to the large design space in our experiment, exhaustive search methods like MoDSE become impractical. Additionally, methods similar to BOOMExplorer can only be applied to the GPR model. Therefore, we select the random sample search method as the baseline for comparison with the ABA algorithm. Fig. 13 visualizes the PHV curves and ADRS curves across iterations, comparing them with the random sample search method.
Figure 13. The convergence of ABA algorithm for IPC-Power optimization.
Regarding PHV convergence, the ABA algorithm notably achieves a higher PHV early on, consistently surpassing the random sample search method throughout various simulation iterations. Specifically, the ABA algorithm reaches a PHV of 38.2 within one hundred iterations, while the random sample search method requires more than 5$\times$ as many iterations to achieve similar levels. This efficiency results from the ABA algorithm’s effective analysis of parameter interactions and their impact on performance metrics, significantly reducing exploration time by 80% and guiding optimal parameter updates for enhanced performance.
Concerning ADRS convergence, it serves to evaluate the quality of estimated Pareto optimal sets in multi-objective optimization by comparing them against the true Pareto set, reflecting the distribution of searched Pareto optimal design points. The ABA algorithm exhibits smoother curves, indicating a more uniform distribution of design points that cover diverse design requirements. In contrast, the random sample search method faces challenges in promptly reaching optimal intervals due to insufficient information on defects in current design points and challenges in selecting subsequent design points based on data distribution.
8. Conclusion
In this paper, we introduced AttentionDSE, a comprehensive DSE framework tailored for high-dimensional design spaces. AttentionDSE seamlessly synergizes the surrogate model with the acquisition function by incorporating the attention mechanism. The innovative PDA mechanism enhances prediction efficiency, while the ABA algorithm offers detailed insights into individual micro-architectural parameter contributions. These advancements improve the design process’s interpretability and minimize the need for expert knowledge. AttentionDSE represents the first attempt to apply attention mechanisms to DSE tasks, showcasing the potential of this approach in optimizing complex design spaces.

### v2

@@
 Multi-objective Optimization in CPU Design Space Exploration: Attention is All You Need
-Abstract.
-Design space exploration (DSE) enables architects to systematically evaluate various design options, guiding decisions on the most suitable configurations to meet specific objectives such as optimizing performance, power, and area.
-However, the growing complexity of modern CPUs has dramatically increased the number of micro-architectural parameters and expanded the overall design space, making DSE more challenging and time-consuming.
-Existing DSE frameworks struggle in large-scale design spaces due to inaccurate models and limited insights into parameter impact, hindering efficient identification of optimal micro-architectures within tight timeframes.
-In this work, we introduce AttentionDSE [^1] The code is now available on anonymous GitHub..
-Its key idea is to use the attention mechanism to establish a direct mapping of micro-architectural parameters to their contributions to predicted performance.
-This approach enhances both the prediction accuracy and interpretability of the performance model.
-Furthermore, the weights are dynamically adjusted, enabling the model to respond to design changes and effectively pinpoint the key micro-architectural parameters/components responsible for performance bottlenecks.
-Thus, AttentionDSE accurately, purposefully, and rapidly discovers optimal designs.
-Experiments on SPEC 2017 demonstrate that AttentionDSE significantly reduces exploration time by over 80% and achieves 3.9% improvement in Pareto Hypervolume compared to state-of-the-art DSE frameworks while maintaining superior prediction accuracy and efficiency with an increasing number of parameters.
-1. Introduction
-Design space exploration (DSE) is an essential phase in CPU design, characterized by an iterative exploration that focuses on balancing performance, power, and area (PPA) (PPA1, ; PPA2, ; PPA3, ; PPA4, ).
-Typically, the DSE can be formulated as a Bayesian optimization (BO) process to find multi-objective trade-offs.
-A prevailing DSE framework primarily comprises two components: the surrogate model and the acquisition function.
-Specifically, the surrogate model predicts the PPA of micro-architectural parameters, bypassing time-consuming simulations (PPA2, ; blackbox2, ; blackbox3, ; BOOMExplorer, ; MoDSE, ), and the acquisition function identifies candidate optimal alternatives for multi-objective optimization, guiding the DSE iterations (nonblackbox1, ; nonblackbox2, ; nonblackbox3, ; nonblackbox4, ; ActBoost, ; bottleneck_analysis1, ).
-Concurrently, the growing complexity of modern CPUs (PowerPC_cpu, ; AMD_cpu, ; superscalar_cpu, ) results in an exponential increase in micro-architectural parameters, known as the high-dimensional design space.
-High-dimensional DSE imposes significant limitations on current DSE frameworks, making it difficult to identify optimal designs within a reasonable timeframe, with some even failing to find optimal solutions altogether.
-Specifically, it presents three key challenges.
+Abstract
+Design Space Exploration (DSE) is essential to modern CPU design, yet current frameworks struggle to scale and generalize in high-dimensional architectural spaces.
+As the dimensionality of design spaces continues to grow, existing DSE frameworks face three fundamental challenges: (1) reduced accuracy and poor scalability of surrogate models in large design spaces; (2) inefficient acquisition guided by hand-crafted heuristics or exhaustive search; (3) limited interpretability, making it hard to pinpoint architectural bottlenecks.
+In this work, we present AttentionDSE, the first end-to-end DSE framework that natively integrates performance prediction and design guidance through an attention-based neural architecture.
+Unlike traditional DSE workflows that separate surrogate modeling from acquisition and rely heavily on hand-crafted heuristics, AttentionDSE establishes a unified, learning-driven optimization loop, in which attention weights serve a dual role: enabling accurate performance estimation and simultaneously exposing the performance bottleneck.
+This paradigm shift elevates attention from a passive representation mechanism to an active, interpretable driver of design decision-making.
+Key innovations include: (1) a Perception-Driven Attention mechanism that exploits architectural hierarchy and locality, scaling attention complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n)$ via sliding windows; (2) an Attention-aware Bottleneck Analysis that automatically surfaces critical parameters for targeted optimization, eliminating the need for domain-specific heuristics.
+Evaluated on high-dimensional CPU design space using the SPEC CPU2017 benchmark suite, AttentionDSE achieves up to 3.9% higher Pareto Hypervolume and over 80% reduction in exploration time compared to state-of-the-art baselines.
+footnotetext: Corresponding author is Mingyu Yan (yanmingyu@ict.ac.cn).
+I Introduction
+As modern CPUs continue to integrate more aggressive micro-architectural features and support increasingly diverse workloads, design space exploration (DSE) has become a crucial phase in the processor design cycle.
+DSE seeks to identify optimal configurations across a vast space of micro-architectural parameters under multiple conflicting objectives, such as performance, power, and area (PPA) [21, 1, 12, 31].
+To accelerate this process, prior work often formulates DSE as a multi-objective Bayesian Optimization (BO) task, leveraging a surrogate model to approximate PPA [1, 6, 9, 3, 42] and an acquisition function to guide design refinement [18, 19, 29, 38, 22, 28].
+However, with the escalating complexity of CPU micro-architecture, the dimensionality of the design space is growing rapidly—reaching hundreds of tunable parameters [26, 27, 20].
+This high-dimensional design space presents three critical challenges for existing DSE frameworks:
   • $\bullet$
-    First, the effectiveness of DSE frameworks is constrained by the limited accuracy and speed of surrogate models in high-dimensional design spaces.
-To expedite DSE, various statistical regression methods (ActBoost, ; MoDSE, ; BO1, ; BO2, ; BOOMExplorer, ; Archranker, ) have been proposed as surrogate models to predict the performance metrics of micro-architecture.
-However, these models suffer from an exponential increase in training and inference time (random_forest, ; GPR, ) and their accuracy decreases with the expanding of the design space (dimgpr, ; dimencurse, ).
+    Inaccuracy and poor scalability of surrogate models: Existing statistical models (e.g., GPR, ensemble trees) struggle to maintain accuracy and scalability in high-dimensional spaces [22, 42, 16, 34, 8, 14], suffering from sparsity of data and a long training time [5, 45].
   • $\bullet$
-    Second, the effectiveness of DSE frameworks is limited by the inefficiency of acquisition functions in high-dimensional design spaces.
-Acquisition functions are crucial in the DSE framework, as they identify the candidate optimal design for exploration.
-However, most acquisition functions depend heavily on expert knowledge to identify the most potential sample in the design space (ActBoost, ; MoDSE, ; BOOMExplorer, ), or on complex feature engineering (ArchExplorer, ) which requires a time-consuming process to identify the bottleneck of the current micro-architecture.
-Moreover, some of them even depend on exhaustive search (MoDSE, ; BOOMExplorer, ; ActBoost, ), which makes the acquisition functions impractical in high-dimensional design spaces.
+    Ineffective acquisition functions: Most acquisition strategies rely on heuristic rules [42, 3], feature engineering [2], or exhaustive search [42, 3, 22]—which become impractical as the micro-architectural parameter count increases.
   • $\bullet$
-    Third, recent DSE frameworks overlook the impact of individual micro-architectural parameters on final performance metrics, missing valuable insights into their contributions.
-This issue primarily arises from the surrogate model’s inability to accurately capture the contribution of each parameter to overall performance (MoDSE, ; BOOMExplorer, ; ActBoost, ; PPA1, ; challenge3_3, ).
-Additionally, the separation of the surrogate model and acquisition function as independent entities in previous efforts further exacerbates this issue, as it impedes the seamless information exchange between them (ArchExplorer, ; MoDSE, ; BOOMExplorer, ; ActBoost, ).
-In this work, we introduce AttentionDSE, a novel DSE framework that leverages the attention mechanism to accurately, purposefully, and rapidly identify optimal CPU designs in the high-dimensional design space.
-The key idea of AttentionDSE lies in its use of attention-based models to establish a direct mapping between micro-architectural parameters and performance metrics.
-This approach improves both the accuracy and interpretability of the performance model.
-Additionally, the attention weights are dynamically adjusted throughout the exploration process, allowing the model to adapt efficiently to design changes.
-This dynamic adjustment enables more effective identification of key micro-architectural parameters responsible for performance bottlenecks.
-As a result, AttentionDSE not only accelerates the design space exploration process but also ensures more precise and purposeful optimization, enabling rapid discovery of optimal micro-architectural designs.
-Overall, our contributions are as follows:
+    Lack of parameter-wise interpretability: Prior work offers limited visibility into the performance sensitivity or bottleneck contribution of each micro-architectural parameters [42, 3, 22], impeding informed design decisions.
+Opportunity.
+We observe that the above challenges are fundamentally rooted in the limited expressiveness and decoupled nature of existing DSE frameworks, where the surrogate model and acquisition function are treated as separate black boxes.
+Meanwhile, recent advances in attention-based deep learning—particularly transformers—offer a compelling alternative.
+First, attention mechanisms excel at modeling long-range dependencies across high-dimensional inputs, enabling them to accurately learn how diverse micro-architectural parameters jointly influence PPA.
+Second, the modular and parallelizable nature of attention layers ensures scalability as the number of micro-architectural parameters increases.
+Third, attention naturally learns to weight the influence of individual parameters when predicting target performance metrics, providing built-in interpretability that is essential for identifying architectural bottlenecks and guiding targeted design refinements.
+These properties make attention-based models uniquely positioned to address the key challenges of DSE.
+By unifying surrogate prediction and acquisition within an attention-driven neural architecture, it becomes possible to achieve accurate, scalable, and explainable exploration—without relying on hand-crafted features or domain-specific rules.
+This convergence of modeling power, transparency, and scalability presents a timely opportunity to rethink DSE methodology for the era of high-dimensional CPU design.
+Our Proposal.
+To address the limitations of existing frameworks in high-dimensional CPU DSE, we propose AttentionDSE—an attention-driven framework that enables end-to-end, interpretable, and scalable DSE.
+AttentionDSE unifies performance prediction and exploration guidance within a single transformer-based neural architecture.
+Its predictor captures both the performance sensitivity and bottleneck contribution of each micro-architectural parameter, with attention weights serving a dual role: effectively supporting accurate PPA estimation and naturally exposing parameter-wise bottlenecks to directly guide the optimization process in a data-driven and interpretable manner.
+To further enhance scalability and interpretability in high-dimensional settings, two key innovations are introduced:
   • $\bullet$
-    We transform the DSE into a sequence prediction problem, introducing AttentionDSE that synergizes the surrogate model and acquisition function and leverages the attention mechanism for high-dimensional DSE on CPU.
+    Perception-Driven Attention (PDA): A hardware-aware serialization and localized attention mechanism that captures architectural hierarchy and local connectivity, reducing attention complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n)$ while greatly improving scalability without loss of modeling fidelity.
   • $\bullet$
-    We introduce a Perception-Driven Attention (PDA) mechanism, which utilizes micro-architectural insights to focus on relevant parameters, minimizing computational overhead of irrelevant parameters to reduce the training and inference time of attention-based performance model.
+    Attention-aware Bottleneck Analysis (ABA): An exploration algorithm that interprets the attention weights to automatically identify performance-critical parameters, enabling fine-grained and explainable design updates without relying on expert knowledge or external heuristics.
+Together, these components enable AttentionDSE to form a closed-loop, data-driven DSE workflow that supports scalable, interpretable, and automated exploration in high-dimensional design spaces.
+This accelerates convergence and yields actionable insights into micro-architectural trade-offs—paving the way for more intelligent and insight-driven architecture design.
+We evaluate AttentionDSE on the SPEC CPU 2017 benchmark suite using an OoO CPU with high-dimensional design space.
+Compared to state-of-the-art DSE frameworks (e.g., BOOMExplorer [3], MoDSE [42], ArchExplorer [2]), AttentionDSE reduces exploration time by over 80%, achieves up to 3.9% improvement in Pareto Hypervolume (PHV), and maintains stable accuracy even in high-dimensional settings.
+Contributions. This work makes the following contributions:
   • $\bullet$
-    We propose an Attention-aware Bottleneck Analysis (ABA) algorithm, which uses attention weights to purposefully and quickly pinpoint bottlenecks, guiding the exploration.
+    We propose AttentionDSE, the first end-to-end attention-driven DSE framework that unifies performance prediction and exploration within a single interpretable architecture.
   • $\bullet$
-    Experimental results show that AttentionDSE reduces exploration time by over 80% and improves Pareto Hypervolume (PHV) by 3.9% compared to state-of-the-art (SOTA) DSE frameworks.
-Additionally, as the number of micro-architectural parameters increases, AttentionDSE keeps stable prediction accuracy and avoids significant increases in training and inference time.
-2. Background
+    We introduce PDA, a hardware-aware serialization and sliding attention mechanism that improves scalability in high-dimensional design spaces.
+  • $\bullet$
+    We develop ABA, a novel method that uses attention weights for interpretable, automated performance bottleneck analysis and design refinement.
+  • $\bullet$
+    We demonstrate significant improvements in accuracy, efficiency, and scalability over state-of-the-art DSE baselines on realistic CPU benchmarks.
+II Background
 In this section, we introduce the relevant concepts.
-2.1. Attention Mechanism
+II-A Attention Mechanism
 The attention mechanism has become one of the most crucial concepts in deep learning.
@@
 A typical attention layer in a neural network dynamically identifies the most relevant parts of the input. The attention score is defined as:
-(1) $\operatorname{Attention}(Q,K,V)=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V.$
+$\operatorname{Attention}(Q,K,V)=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V.$ (1)
 It operates using three key vectors: query (Q), key (K), and value (V).
@@
 This process is especially effective in self-attention, where each input element attends to all others, and in multi-head attention, which captures different aspects of the input through multiple attention heads.
-Attention layers are foundational in models like transformers (transformer, ), enabling flexible and efficient learning of dependencies within sequences.
-2.2. Design Space Exploration in CPU Design
-DSE is a critical technique in the optimization and design process, primarily used to identify and analyze various design options within CPU design (PPA1, ; PPA2, ; PPA3, ; PPA4, ).
-Its goal is to systematically explore all possible design alternatives to find the optimal or near-optimal solutions that meet specific design requirements and constraints (ArchExplorer, ; MoDSE, ; BOOMExplorer, ; ActBoost, ; wangduo1, ; wangduo2, ).
-To expedite the process, prevailing DSE frameworks adhere to the fundamental principles of BO (MoDSE, ; BOOMExplorer, ; ActBoost, ; Archranker, ).
-BO leverages a surrogate model that serves as a predictive tool, approximating the underlying objective function (GPR, ; random_forest, ; ensemble_learning, ).
-Additionally, an acquisition function drives the optimization process in DSE frameworks, efficiently navigating the DSE by identifying promising design points (PPA2, ), which is a set of micro-architectural parameters, for evaluation (ArchExplorer, ).
-The above two facilitate effective exploration of the design space within a reasonable time (BO1, ; BO2, ).
-Figure 1. The workflow of prevailing DSE framework. ❶~❹ form the exploration loop.
+Attention layers are foundational in models like transformers [40], enabling flexible and efficient learning of dependencies within sequences.
+II-B Design Space Exploration in CPU Design
+DSE is a critical technique in the optimization and design process, primarily used to identify and analyze various design options within CPU design [21, 1, 12, 31].
+Its goal is to systematically explore all possible design alternatives to find the optimal or near-optimal solutions that meet specific design requirements and constraints [2, 42, 3, 22, 43, 41].
+To expedite the process, prevailing DSE frameworks adhere to the fundamental principles of BO [42, 3, 22, 10].
+BO leverages a surrogate model that serves as a predictive tool, approximating the underlying objective function [45, 5, 13].
+Additionally, an acquisition function drives the optimization process in DSE frameworks, efficiently navigating the DSE by identifying promising design points [1], which is a set of micro-architectural parameters, for evaluation [2].
+The above two facilitate effective exploration of the design space within a reasonable time [16, 34].
+Figure 1: The prevailing DSE workflow. ❶˜❹ form the exploration loop.
 Fig.
 1 illustrates the core workflow of the BO-based DSE framework.
-Prior to the exploration stage, the framework first trains the surrogate model using sampling methods, such as Plackett-Burman ranking(PBrank, ; MoDSE, ), to generate the training set (①, ②, ③).
+Prior to the exploration stage, the framework first trains the surrogate model using sampling methods, such as Plackett-Burman ranking[39, 42], to generate the training set (①, ②, ③).
 Once the surrogate models are trained, the exploration stage begins.
@@
 The framework then iteratively updates the Pareto-optimal set and repeats this process (❸, ❹).
-2.3. Pareto Optimal Set
+II-C Pareto Optimal Set
 A key concept in CPU DSE is the Pareto optimal set.
@@
 In an n-objective minimization problem, a design point $x$ is defined to be dominated by $x^{*}$ if
-(2) $\forall i\in[1,n],f_{i}\left(x^{*}\right)\leq f_{i}(x);\exists j\in[1,n],f_{j}\left(x^{*}\right)<f_{j}(x).$
+$\forall i\in[1,n],f_{i}\left(x^{*}\right)\leq f_{i}(x);\exists j\in[1,n],f_{j}\left(x^{*}\right)<f_{j}(x).$ (2)
 This dominance is denoted as the partial order $x^{*}\prec x$ and $y^{*}\prec y$.
@@
 Among all design points, the set of design points that are not dominated by any other design points is called the Pareto optimal set $\Omega$, formulated as:
-(3) $\Omega=\left\{x\mid x^{*}\nprec x,\forall x^{*}\in D\right\}.$
+$\Omega=\left\{x\mid x^{*}\nprec x,\forall x^{*}\in D\right\}.$ (3)
 The PHV is determined by the Pareto optimal set and a reference point, as shown in Fig.2(a).
@@
 In Fig.2(b), the new design points $x$ and $y$ demonstrate superior performance or power compared to the original design points, with both expanding the PHV, illustrated by the red and purple regions, respectively.
-Figure 2. The Expansion of PHV: (a) original PHV; (b) design points excelling in performance or power extend the PHV.
-3. Related Works
+Figure 2: The improvement of PHV: (a) original PHV; (b) design points excelling in performance or power extend the PHV.
+III Related Works
+Many efforts [10, 35, 44, 24, 3, 42, 2, 25, 22] accomplish the DSE tasks.
 For the early work, Mariani et al.
-(relative_work1, ) use a Kriging model (Kriging, ) with exponential correlation as a surrogate to predict the Pareto nondominated rank distribution.
+[24] use a Kriging model [32] with exponential correlation as a surrogate to predict the Pareto nondominated rank distribution.
 They then iteratively optimize using the Expected Improvement (EI) of non-dominated levels as the acquisition function.
-CASH (relative_work2, ) compares multiple machine learning models and finds that BO achieves superior exploration results, by using a random forest as the surrogate model and EI as the acquisition function.
+CASH [35] compares multiple machine learning models and finds that BO achieves superior exploration results, by using a random forest as the surrogate model and EI as the acquisition function.
 Wang et al.
-(relative_work3, ) employ an adaptive component selection and smoothing operator as the surrogate model and the expected Hypervolume Improvement (HVI) of the Pareto frontier as the acquisition function.
-For the blackbox DSE framework, ActBoost (ActBoost, ) integrates statistical methods and machine learning to prune the design space and determine the training set.
-For the surrogate model, ActBoost uses two AdaBoost.RT (AdaBoost.RT, ) models with active learning.
+[44] employ an adaptive component selection and smoothing operator as the surrogate model and the expected Hypervolume Improvement (HVI) of the Pareto frontier as the acquisition function.
+For the blackbox DSE framework, ActBoost [22] integrates statistical methods and machine learning to prune the design space and determine the training set.
+For the surrogate model, ActBoost uses two AdaBoost.RT [36] models with active learning.
 For the acquisition function, it calculates the prediction uncertainty of each unsampled design point and selects one with the highest uncertainty.
 By combining statistical sampling and the boosting technique, ActBoost enables efficient and accurate DSE tasks.
-BOOMExplorer (BOOMExplorer, ) employs Gaussian Process Regression (GPR) (GPR, ) as the surrogate model and leverages EI (EI, ) as the acquisition function, searching for the next design point by analyzing data distribution characteristics.
+BOOMExplorer [3] employs Gaussian Process Regression (GPR) [45] as the surrogate model and leverages EI [46] as the acquisition function, searching for the next design point by analyzing data distribution characteristics.
 By embedding prior knowledge, BOOMExplorer achieves superior compared to analytical methods.
-MoDSE (MoDSE, ) leverages ensemble learning for more accurate prediction.
-It combines AdaBoost (AdaBoost, ) and Gradient Boosted Regression Trees (GBRT) (GBRT, ) as the surrogate model and uses HVI with a full-space search for the acquisition function.
+MoDSE [42] leverages ensemble learning for more accurate prediction.
+It combines AdaBoost [33] and Gradient Boosted Regression Trees (GBRT) [30] as the surrogate model and uses HVI with a full-space search for the acquisition function.
 MoDSE introduces a uniformity-aware selection algorithm for efficient multi-objective DSE and a Pareto-rank-based sample weight generation algorithm to train the surrogate model, enhancing the DSE outcomes.
-For the DSE frameworks based on bottleneck analysis, ArchExplorer (ArchExplorer, ) adopts the dynamic event-dependence graph (DEG) (DEG, ) for critical path analysis and identifies bottlenecks in the current micro-architecture to guide the design of micro-architecture.
-It couples this approach with a software simulator (Gem5) (gem5, ) as the surrogate model.
+For the DSE frameworks based on bottleneck analysis, ArchExplorer [2] adopts the dynamic event-dependence graph (DEG) [15] for critical path analysis and identifies bottlenecks in the current micro-architecture to guide the design of micro-architecture.
+It couples this approach with a software simulator (Gem5) [4] as the surrogate model.
 ArchExplorer focuses on refining the acquisition function to reduce the need for domain knowledge in mechanistic models.
-Explainable-DSE (Explainable_DSE, ) is the SOTA framework for DSE tasks in deep neural network (DNN) accelerators.
+Explainable-DSE [11] is the SOTA framework for DSE tasks in deep neural network (DNN) accelerators.
 By leveraging the structure and characteristics of neural networks, Explainable-DSE constructs a bottleneck model to identify design bottlenecks.
@@
 This framework is primarily used for the joint DSE of DNN accelerators and DNN algorithms.
-4. Motivation
+IV Motivation
 In this section, we will explain the motivations behind leveraging the attention mechanism in the DSE framework.
-4.1. Challenges in Current DSE Studies
-Modern CPU design is becoming increasingly complex, with a growing number of micro-architectural parameters.
-The exploration in such a significantly large design space is known as the high-dimensional DSE.
-Current DSE frameworks face three primary challenges in high-dimensional tasks.
-First, DSE frameworks struggle with effectiveness due to the limited accuracy and efficiency of surrogate models in high-dimensional design spaces.
-The surrogate models in prevailing DSE frameworks (ActBoost, ; MoDSE, ; BOOMExplorer, ; Archranker, ), which typically use statistical regression methods, are simple while vulnerable.
-Specifically, prediction accuracy can decline as dimensionality increases.
-For instance, when the number of parameters exceeds 75, BOOMExplorer struggles to accurately predict and associate micro-architectural parameters with performance metrics.
-This issue arises from the sparser feature space in high-dimensional condition (dimencurse, ), indicating the need for more training data to accurately capture the data distribution.
-However, the training and inference time increases greatly as the size of the training set grows.
-For instance, with a training set of size $n$, the complexity of training the inverse of the covariance matrix typically reaches $O(n^{3})$ and the inference through the inverse of the covariance matrix is $O(n^{2})$.
-This considerable computational cost renders high-dimensional design spaces impractical.
-Second, the performance of DSE frameworks is hindered by the inefficiency of acquisition functions when dealing with high-dimensional design spaces.
-Acquisition functions are crucial for finding optimal design configurations, but they face challenges in high-dimensional design spaces.
-Current methods often rely on expert knowledge (ActBoost, ; MoDSE, ; BOOMExplorer, ) or complex feature engineering (ArchExplorer, ) to identify the bottleneck of the current micro-architecture, which can introduce biases, require significant time, and lack scalability.
-For instance, ArchExplorer only analyses the first hundred thousand of the instructions and achieves the optimal architecture with the number of parameters 21 over 15 days (ArchExplorer, ).
-Moreover, some methods rely on large-scale or even exhaustive search (MoDSE, ; ActBoost, ; BOOMExplorer, ).
-In high-dimensional spaces, this exhaustive search becomes computationally prohibitive due to the enlarged design space.
-For example, MoDSE can only perform with the number of parameters 10 (totaling 36,864 design points) (MoDSE, ).
-When the number of parameters increases to 30, it may take $1.9\times 10^{29}$ years.
-The high computational cost of such searches slows down the DSE process, delaying the identification of high-performance designs.
-This inefficiency slows down the DSE process, highlighting the need for more advanced and efficient acquisition strategies to handle complex, high-dimensional problems.
-Third, conventional DSE frameworks fail to recognize the impact of each parameter on performance metrics.
-This limitation can cause these frameworks to overlook critical insights into how individual parameters impact the performance metrics, making it difficult to quickly and accurately identify bottlenecks in the current design.
-This phenomenon can be attributed to two primary reasons.
-Firstly, existing surrogate models cannot unveil the contributions of each parameter to the performance metrics because they rely on mathematical analysis (ActBoost, ; Archranker, ) or simplistic black-box models (BOOMExplorer, ; MoDSE, ) for prediction.
-Secondly, prevailing DSE frameworks typically adhere to a two-stage DSE scheme, treating the surrogate model and acquisition function as independent entities.
-This separation inhibits the exchange of information, preventing a comprehensive understanding of parameter efficacy (BOOMExplorer, ; MoDSE, ; ActBoost, ; ArchExplorer, ).
-Moreover, the lack of integration between these two components hinders the ability of the acquisition function to adaptively refine the exploration results based on feedback from the surrogate model, resulting in suboptimal performance and less effective DSE.
-4.2. Opportunity of Attention in DSE Tasks
-DSE is a crucial process in CPU design, often consisting of two important tasks: prediction and bottleneck analysis.
-Prediction helps estimate the performance of various design configurations, while bottleneck analysis identifies critical constraints that limit system efficiency.
-Both tasks are essential to guide CPU design effectively.
-Recent advances in attention mechanisms offer significant advantages in addressing these challenges, proving to be highly effective for both performance prediction and bottleneck analysis.
-First, attention has been demonstrated to excel in regression tasks (attention_regression1, ; attention_regression2, ), accurately capturing complex relationships between inputs and outputs.
-Second, attention mechanisms inherently highlight the importance of different parameters by assigning varying weights, thus enabling a more fine-grained analysis of key design factors.
-These strengths align directly with the core needs of DSE, where both accurate performance prediction and the identification of critical design parameters are essential for driving effective design exploration.
-5. Design of AttentionDSE
-In this section, we introduce AttentionDSE, the first DSE framework that adopts the attention mechanism, suitable for high-dimensional DSE.
-5.1. Overview of AttentionDSE
-Figure 3 provides an overview of AttentionDSE.
-It comprises three primary components: 1) Predictors, which provide rapid and accurate predictions of performance metrics (e.g., IPC) for given micro-architectures across benchmarks, thereby obviating the need for time-intensive simulations; 2) Micro-architecture Serialization, designed to reduce the computational burden associated with predictions, thereby expediting the performance evaluation process; 3) Bottleneck Analysis and Iterator, which identify the micro-architectural parameters that introduce performance bottlenecks within micro-architecture modules, facilitating the exploration of promising micro-architectural candidates by selectively modifying these parameters.
+IV-A Challenges in Current DSE Studies
+As modern CPUs continue to integrate increasingly aggressive micro-architectural features and support a wider range of workloads, their design complexity grows substantially.
+This complexity is reflected in the expanding number of tunable micro-architectural parameters—spanning pipeline width, core frequency, and so on.
+As a result, the design space has become significantly larger and more intricate, giving rise to the challenge of high-dimensional DSE.
+Existing DSE frameworks, however, struggle to scale in such settings and face three fundamental challenges in high-dimensional exploration tasks.
+First, the effectiveness of DSE frameworks is limited by the low accuracy and poor scalability of surrogate models in high-dimensional CPU design spaces.
+Most existing approaches [22, 42, 3, 10] use statistical models—such as linear regression or Gaussian processes—to predict PPA.
+While efficient on small search spaces, these models quickly break down as the number of design parameters grows.
+For instance, when the design space exceeds 75 parameters, BOOMExplorer struggles to maintain prediction accuracy.
+This is mainly due to the “curse of dimensionality”: as the number of parameters increases, the design space becomes exponentially sparser, requiring significantly more samples to train reliable models [14].
+Unfortunately, increasing training size leads to excessive overhead.
+For example, in Gaussian Process-based models, training requires matrix inversion with $O(n^{3})$ complexity and inference scales as $O(n^{2})$, where $n$ is the number of training samples.
+This makes it impractical to scale traditional surrogate modeling to realistic, high-dimensional CPU design scenarios.
+Second, the performance of DSE frameworks is hindered by the inefficiency of acquisition functions in high-dimensional design spaces.
+Acquisition functions are critical for guiding exploration toward optimal micro-architectural configurations.
+However, in large parameter spaces, existing strategies become inefficient and poorly scalable.
+Many methods depend on domain-specific heuristics [22, 42, 3] or handcrafted features [2] to identify performance bottlenecks.
+These approaches are often biased, time-consuming, and difficult to generalize across architectures and workloads.
+For example, ArchExplorer [2] analyzes only the first 100,000 instructions and takes 15 days to optimize an architecture with just 21 parameters.
+Some frameworks resort to large-scale or even exhaustive search [42, 22], which becomes computationally infeasible as dimensionality increases.
+MoDSE [42], for instance, operates over a 10-parameter space with 36,864 design points, but scaling to 30 parameters would require an astronomical $1.9\times 10^{29}$ years of search time.
+These limitations highlight the need for acquisition strategies that can operate efficiently in high-dimensional settings—without relying on heuristics or exhaustive enumeration—while still identifying high-quality designs.
+Third, existing DSE frameworks lack interpretability, making it difficult to reason about trade-offs and locate architectural bottlenecks.
+Conventional frameworks offer little insight into how individual micro-architectural parameters affect PPA.
+This makes it challenging to understand why a given design performs well or poorly, thereby hampering informed, targeted design refinements.
+This problem stems from two core limitations.
+First, most surrogate models rely on either mathematical analysis [22, 10] or black-box predictors [3, 42], which fail to expose parameter-wise contributions to performance.
+Second, the standard two-stage DSE workflow treats performance prediction and design exploration as separate steps [3, 42, 22, 2], preventing feedback from the predictor from guiding future exploration.
+As a result, it is difficult for the acquisition function to adaptively prioritize parameters that are more performance-sensitive.
+Without a unified, feedback-driven approach, existing methods are not only harder to interpret, but also less effective at navigating high-dimensional design spaces-especially when timely, bottleneck-aware decisions are required.
+IV-B Why Attention for DSE?
+The attention mechanism provides a powerful and well-suited solution to above challenges for several reasons:
+  • $\bullet$
+    Context-aware Interaction Modeling: Attention dynamically computes the relative importance of each parameter with respect to others in a given design configuration [7].
+This enables the model to capture complex, conditional dependencies—such as how the benefit of a wide fetch stage may depend on the reorder buffer size or instruction window depth—without manual feature engineering.
+  • $\bullet$
+    Scalability and Structural Flexibility: With recent advances such as sparse attention, attention-based models can scale to hundreds of parameters while maintaining modeling fidelity.
+Moreover, attention can be adapted to reflect architectural hierarchies and locality, enabling better alignment with the structure of CPU design spaces.
+  • $\bullet$
+    Unified Modeling for Prediction and Guidance: Attention can bridge surrogate modeling and exploration by using a single predictor to both estimate PPA and inform the search direction.
+This eliminates the need for hand-crafted heuristics or separate analysis stages and enabling a more efficient, closed-loop DSE process.
+  • $\bullet$
+    Interpretability through Attention Weights: Unlike black-box predictors, attention mechanisms expose internal weights that reveal which parameters contribute most to performance metrics.
+This provides a natural interface for bottleneck analysis and informed refinement, supporting more targeted and efficient exploration.
+These properties position attention not merely as a prediction tool, but as a unified interface between surrogate modeling and acquisition strategy.
+When properly integrated, attention mechanisms can enable an end-to-end, self-driven DSE framework that is both interpretable and scalable—paving the way for a new class of intelligent architecture design tools.
+V Design of AttentionDSE
+In this section, we introduce AttentionDSE, an attention-based framework designed for high-dimensional DSE.
+V-A Overview of AttentionDSE
+Fig.
+3 provides an overview of AttentionDSE.
+It comprises three primary components: (1) Predictors, which provide rapid and accurate predictions of performance metrics (e.g., IPC) for given micro-architectures across benchmarks; (2) Micro-architecture Serialization, designed to convert various design parameters into input vectors for the predictor while ensuring the positional certainty of parameters; it also reduces the computational burden associated with predictions, thereby expediting the performance evaluation process; (3) Bottleneck Analysis and Iterator, which identify the micro-architectural parameters that introduce performance bottlenecks within micro-architecture modules, facilitating the exploration of promising micro-architectural candidates by selectively modifying these parameters.
 The process of the AttentionDSE consists of a predictor training stage (①, ②, ③, ④) and an architecture exploration stage (❶, ❷, ❸, ❹, ❺).
-Figure 3. Overview of the AttentionDSE.
-Figure 4. The workflow of AttentionDSE framework in the exploration stage.
+Figure 3: Overview of the AttentionDSE.
+Figure 4: The workflow of AttentionDSE framework in the exploration stage.
 The training stage is a one-time process, and the sequence of operations is represented by hollow numbers.
@@
 Next, the deep learning model, integrated with the PDA, functioning as the predictor, is trained to predict performance metrics (④).
-Benefiting from probabilistic attribute correlation in attention calculation, it’s more convenient to capture the intricate relationship between parameters, which informs exploration hints for the exploration stage.
+Benefiting from probabilistic attribute correlation in attention calculation, it more easily captures the intricate relationship between parameters, informing hints for the exploration stage.
 Once the predictors are trained, AttentionDSE is ready for the exploration stage, which is iterative and aims for efficient multi-objective optimization.
@@
 All these techniques will be discussed in detail in the following sections.
-5.2. Attention-based Prediction Model
-A fast and accurate performance predictor can replace traditional simulation methods, serving as the cornerstone of an efficient DSE framework.
-This approach significantly accelerates design iterations and reduces the overall design cycle.
-As discussed in Section 4, the attention-based model is a promising option for developing such a predictor.
-Neural networks possess strong nonlinear fitting capabilities, making them well-suited for learning from high-dimensional data.
-Additionally, they benefit from robust hardware acceleration and a well-developed GPU-based framework, collectively enhancing training and inference speed.
-In this subsection, we propose an attention-based prediction model as the single-objective predictor.
-We aim to develop a more accurate and scalable predictor that emphasizes efficiency in high-dimensional DSE.
-Input: $D$: Design space; $n$: The size of the training set; $k$: Embedding length; $Depth$: The depth of the model; $Epoch$: The total epochs for training.
+V-B Attention-based Prediction Model
+A fast and accurate performance predictor is essential to replacing traditional simulation-based methods, forming the foundation of an efficient DSE framework.
+Such predictors enable rapid estimation of performance metrics, significantly accelerating design iterations and reducing overall exploration time.
+As discussed in Section IV, attention-based neural networks are particularly well-suited for this task: they not only excel at modeling complex, context-dependent relationships in high-dimensional design spaces, but also naturally support parallel computation, making them highly compatible with GPU-based acceleration.
+Thus, we introduce an attention-based model tailored for single-objective performance prediction.
+By combining strong modeling capacity with scalable inference, the proposed surrogate supports fast and accurate exploration across large micro-architectural design spaces.
+Input: $D$: Design space; $n$: The size of the training set; $k$: Embedding length; DepthDepth: The depth of the model; EpochEpoch: The total epochs for training.
 Output: $M$: The trained predictor.
@@
 […]
-3 Instantiate the prediction token $P_{t}$ with embedding length $k$;
+3 Instantiate the prediction token $T$ with embedding length $k$;
 4 Initialize the weights of the Predictor;
 5 Simulate $T$ to obtain $S_{ground\_truth}$ of IPC, Power, and Area;
-6 for $epoch\leftarrow 0$ to $Epoch$ do
-7       $S\leftarrow Serialize(T)$;
-8       $S_{embed}\leftarrow Embedding(S)$;
-9       $S_{embed}\leftarrow Concat(S_{embed},P_{t})$;
-10       for $depth\leftarrow 0$ to $Depth$ do
+6 for $epoch\leftarrow 0$ to EpochEpoch do
+7 $S_{embed}\leftarrow Embedding(T)$;
+8 $S_{embed}\leftarrow Add(S_{embed},PE)$;
+9 for $depth\leftarrow 0$ to DepthDepth do
 10    $S_{attn}\leftarrow SelfAttention(S_{embed})$;
 11    $S_{mlp}\leftarrow MLP(S_{attn})$;
-13
+12
 13   end for
-15      $S_{attn},A_{score}\leftarrow SelfAttention(S_{mlp})$;
-16       Extract processed prediction token $S_{attn}\leftarrow S_{attn}[0,:]$;
-17       $S_{pred}\leftarrow FullyConnectedLayer(S_{attn})$;
+14 $P_{t},Attn\_{score}\leftarrow SelfAttention(S_{mlp})$;
+15   Extract analysis token $A_{t}\leftarrow Reduce\_{sum}(Attn\_{score})$;
+16 $S_{pred}\leftarrow FullyConnectedLayer(P_{t})$;
 17 $BackwardPropagation(S_{pred},S_{ground\_truth})$;
-19
+18
 19 end for
-21return Predictor;
+20return Predictor;
 Algorithm 1 The Training Procedure of Predictors
-The training procedure of the attention-based prediction model is demonstrated in Algorithm.
-1.
-The prediction model consists of two key components: embedding preprocessing and the attention-based model.
-For embedding preprocessing, the procedure involves three steps: (a) Embedding, which transforms the discrete and heterogeneous alternatives of each micro-architectural parameter into a continuous numerical representation, simplifying optimization and enhancing the representation of relationships within the micro-architecture; (b) Position Embedding (PE), which assigns positional encoding to each item in the parameter sequence, enabling the model to learn parameter attributes more effectively and constrain the embedding space of each parameter; and (c) Prediction Token (PT), which gathers information from parameter embeddings using the self-attention mechanism and fuses it through multilayer perceptron (MLP), allowing the model to perceive the entire parameter sequence effectively.
-As for the attention-based model, to ensure the generalizability and versatility of the model, the predictor includes only vanilla self-attention, MLP, and normalization operations, which are the minimal necessary components of the transformer model (transformer, ).
-This streamlined design aims to maintain the attention mechanism’s inherent advantages.
-By focusing on these core components, we can leverage the self-attention mechanism’s ability to capture information within the input design point, allowing for the effective handling of complicated relationships between the design point and performance metrics in high-dimensional DSE tasks.
-The inclusion of MLP layers provides the model with the capacity to learn complex patterns and representations, while normalization operations ensure stable and efficient training by mitigating issues related to internal covariate shifts.
-5.3. Perception-driven Attention Mechanism
+Algorithm 1 outlines the training procedure of our attention-based predictor.
+The process begins by randomly sampling $n$ design points from the design space $D$ and simulating each to obtain ground truth metrics (IPC, Power, and Area).
+This procedure comprises the following steps.
+  • $\bullet$
+    Embedding: Each discrete, heterogeneous microarchitectural parameter is mapped to a vector representation, enabling the model to reason over diverse parameter types in a unified latent space.
+  • $\bullet$
+    Position Embedding (PE): Positional encodings are added to retain the identity and ordering of each parameter, allowing the attention mechanism to associate learned importance scores with specific microarchitectural features.
+  • $\bullet$
+    Prediction Token ($P_{t}$): Prediction token aggregates global information via self-attention, and is processed by an MLP to produce final PPA predictions.
+  • $\bullet$
+    Analysis Token ($A_{t}$): Derived from attention scores from the final self-attention layer, $A_{t}$ captures the relative importance of each parameter, enabling bottleneck analysis and interpretability.
+By embedding structural context and aligning parameter positions, the model effectively captures complex interdependencies among architectural parameters—addressing the key challenge of scalability and insight in high-dimensional DSE.
+V-C Perception-driven Attention Mechanism
 Although the prediction time is significantly reduced compared to simulation, further reductions are still necessary, especially when handling the large-scale design spaces of CPUs.
@@
 The PDA mechanism includes two aspects: the micro-architecture serialization method and the sliding window attention mechanism.
-In previous work (GPR, ; MoDSE, ; ensemble_learning, ; BOOMExplorer, ), the order of parameters is not a concern.
-Because GPR (GPR, ; BOOMExplorer, ) relies on the distance and similarity between input data rather than their arrangement order in statistical regression methods.
-Similarly, in ensemble learning (ensemble_learning, ; MoDSE, ), the training process of multiple weak learners introduces randomness, making the order of parameters irrelevant.
-To address this challenge in the attention-based predictor, we introduce a serialization approach named Perception-driven Serialization (PDS), aimed at preserving architectural information and laying the groundwork for the sliding window attention technique.
-This technique employs a graph-based method to delineate relationships among all parameters within each pipeline stage.
-Figure 5. The Perception-driven Serialization approach. (a) A Sample OoO CPU core; (b) Rename Stage; (c) Stage abstraction; (d) The steps of Perception-driven Serialization approach.
+In previous work [45, 42, 13, 3], the order of parameters is not a concern.
+Because GPR [45, 3] relies on the distance and similarity between input data rather than their arrangement order in statistical regression methods.
+Similarly, in ensemble learning [13, 42], the training process of multiple weak learners introduces randomness, making the order of parameters irrelevant.
+To address this challenge in the attention-based predictor, we introduce a serialization approach termed Perception-driven Serialization (PDS), which preserves architectural information and establishes a foundation for the sliding window attention technique.
+This method leverages a graph-based framework to map relationships among parameters within each pipeline stage, rooted in a key insight: CPU instruction and data flows are processed sequentially across pipeline stages, with parameters within the same stage interacting far more frequently than those across different stages.
+Figure 5: The Perception-driven Serialization approach. (a) A Sample OoO CPU core; (b) Rename Stage; (c) Stage abstraction; (d) The steps of Perception-driven Serialization approach.
 An example is illustrated in Fig.
@@
 We chose the Rename stage, shown in Fig.
-5(b), as an example, where there are four parameters to be serialized including the free list (F), the rename map (R), the scoreboard (S), and the history buffer (H).
+5(b), as an example, where there are four parameters to be serialized, including the free list (F), the rename map (R), the scoreboard (S), and the history buffer (H).
 The PDS first transfers the micro-architecture into a perceptual graph, which is an undirected graph based on the dataflow among these components, as depicted in Fig.
@@
 To quantify the perception of each component, we leverage the perception degree $D$ of each parameter in the perceptual graph.
-The perception degree is calculated as follows:
-(4) $D(\text{parameter})=\sum_{e\in E_{\text{parameter}}^{\text{internal}}}e-\sum_{e\in E_{\text{parameter}}^{\text{external}}}e$
+$D$ is calculated as follows:
+$D(\text{parameter})=\sum_{e\in E_{\text{parameter}}^{\text{internal}}}e-\sum_{e\in E_{\text{parameter}}^{\text{external}}}e$ (4)
 Here, the internal edges represent the intra-stage perception field, while the external edges represent the inter-stage perception field.
@@
 After all the pipeline stages have been serialized, the final vector of the micro-architecture is combined according to the order of the pipeline stages.
-Figure 6. The attention weight calculation. (a) Vanilla self-attention mechanism; (b) PDA mechanism.
-After the micro-architecture has been serialized, the vanilla self-attention mechanism can be replaced by the sliding window attention mechanism due to the parameter clusters.
+Figure 6: The attention weight calculation. (a) Vanilla self-attention mechanism; (b) PDA mechanism.
+In the clustered sequence, the correlation between adjacent parameters is significantly higher than that between non-adjacent parameters.
+At this point, the sliding window only needs to cover a local area to capture most of the key dependencies, without the need to calculate global attention.
 Fig.
@@
 In this way, for a parameter vector of length $n$, the computational complexity is reduced from $O(n^{2})$ to $O(n)$.
-Thus, the PDA approach alleviates the limitations posed by number of parameters, enhances scalability, and reduces the training and inference time of the prediction model.
-5.4. Attention-aware Bottleneck Analysis
-Bottleneck analysis is a critical step in CPU design, enabling architects to identify the key factors that impact system performance.
-In this subsection, we propose the ABA algorithm, which equips AttentionDSE with the capability to analyze attention weight heatmaps generated by an attention-based predictor.
-The ABA algorithm, detailed in Algorithm 2, reveals the intricate relationships between micro-architectural parameters and performance metrics, thus guiding the exploration of design space.
-This approach ensures that the most influential design parameters are prioritized during optimization, leading to more efficient and effective DSE.
+Thus, the PDA approach alleviates the limitations posed by the number of parameters, enhances scalability, and reduces the training and inference time of the prediction model.
+V-D Attention-aware Bottleneck Analysis
+Bottleneck analysis is a crucial step in CPU design, enabling architects to identify the key microarchitectural factors that significantly influence performance.
+To support this capability, we propose the ABA algorithm (Algorithm 2), which captures intricate relationships between microarchitectural parameters and performance metrics.
+By identifying the parameters that contribute most to performance variation, ABA effectively guides DSE by prioritizing high-impact parameters, thereby improving both exploration efficiency and design quality.
+To enable such analysis in an interpretable and quantifiable manner, we extract a dedicated analysis token $A_{t}$ from the previously introduced prediction model.
+This token is designed to assess the importance of each input parameter with respect to performance outcomes.
+A central requirement for ensuring the reliability of this mechanism is maintaining a strict one-to-one correspondence between attention weights and the input parameters.
+To ensure this, the model preserves token position consistency across all components: an embedding layer, PDA layer, MLP, self-attention, and fully connected (FC) layer.
+Specifically, the embedding layer applies independent projections to each parameter token, the PDA layer adds positional encoding while preserving order, and both the MLP and self-attention layers operate token-wise to retain the direct mapping between tokens and their associated attention weights.
+This rigorous preservation of token alignment is vital: it enables each attention weight to be directly and explicitly mapped back to a specific design parameter.
+Leveraging this property, the ABA algorithm interprets the attention-weight heatmap generated by the model to identify performance-critical parameters.
+In this heatmap, each row indicates how much a particular parameter is influenced by others, while each column—indexed by the analysis token $A_{t}$—reflects the degree to which a parameter affects all other parameters.
+The column-wise sum of attention weights thus serves as a proxy for a parameter’s global impact on performance prediction.
+Based on this insight, ABA detects bottleneck parameters by analyzing $A_{t}$: a low value for IPC (indicating high sensitivity) or a high value for Power or Area (indicating dominant cost contribution) signifies that the corresponding parameter is a bottleneck for that specific objective.
+ABA then updates the design configuration accordingly, refining the Pareto frontier and facilitating iterative improvements in both design quality and exploration efficiency.
 Input: $D$: The design space; $X$, $X_{n}$: Initial design point set and its size; $I_{max}$: The max iteration of exploration;
@@
 3 $\Omega\leftarrow GetParetoOptimal(Predictor(X))$;
-4 for $i$leftarrow 0 to $I_{max}$ do
+4 for $i$ $\leftarrow$ 0 to $I_{max}$ do
 5   Set search queue $Q\leftarrow\varnothing$;
 6 for each design point x in $\Omega$ do
-7             $x^{\prime}\leftarrow BottleneckAnalysis(x)$;
+7    $A_{t}\leftarrow Predictor(x)$;
+8    $x^{\prime}\leftarrow BottleneckAnalysis(x,A_{t})$;
 9    $PPA\leftarrow Predictor(x^{\prime})$;
@@
 11       $Q.push(x^{\prime})$;
-11
+12
 13      end if
-13
+14
 15   end for
 16 $\Omega\leftarrow GetParetoOptimal(Predictor(Q))$;
-16
+17
 18 end for
-18return $\Omega$;
+19return $\Omega$;
 Algorithm 2 Attention-aware Bottleneck Analysis
-In the attention weight heatmap, each row represents how the current parameter is influenced by other parameters, reflecting the correlations between them.
-For a particular parameter, the sum of all weights in its corresponding column indicates the overall influence of this parameter on others.
-Therefore, if the total sum of attention weights in the column corresponding to a certain parameter is minimal for IPC or maximal for Power or Area, we infer that this parameter has a significant impact on the current performance metrics.
-ABA algorithm then updates the Pareto optimal set by updating the design point based on this bottleneck analysis, ensuring continuous improvement and exploration of the design space.
-Figure 7. The example of ABA algorithm updating the micro-architectural parameters.
+Figure 7: The example of ABA algorithm updating the parameters.
 To clarify the execution process of the ABA algorithm, Fig.
@@
 By strategically adjusting the instruction queue size, the pipeline is provided with a sufficient flow of instructions, further optimizing performance.
-6. Experimental Setup
+VI Experimental Setup
 In this section, we introduce the experimental setup used to assess the performance of AttentionDSE.
 Simulation.
-We utilize GEM5 (gem5, ) as the timing-accurate simulator and McPAT (mcpat, ) as the power and area modeling tool.
+We utilize GEM5 [4] as the timing-accurate simulator and McPAT [23] as the power and area modeling tool.
+McPAT is configured with technology parameters aligned with industry-standard VLSI design rules, enabling it to estimate power (dynamic and leakage) and area based on microarchitectural specifications, which is consistent with prior DSE works [42, 3].
 We extend GEM5 using Python to generate different CPU cores with varying configurations.
-To bolster the credibility of our experiments, we employ SPEC CPU2017 (SPEC2017, ) as the evaluation benchmark.
-For identifying resource utilization status, we utilize Simpoints (simpoints, ) for each workload evaluation.
+To bolster the credibility of our experiments, we employ SPEC CPU2017 [37] as the evaluation benchmark.
+For identifying resource utilization status, we utilize Simpoints [17] for each workload evaluation.
 Each workload is divided into at most 30 clusters, with each cluster containing ten million instructions.
 Baselines.
-Our baselines include the ActBoost (ActBoost, ), the BOOMExplorer (BOOMExplorer, ), and the MoDSE (MoDSE, ).
-Additionally, we implement ArchExplorer (ArchExplorer, ) for comparison with our ABA algorithm.
+Our baselines include the ActBoost [22], the BOOMExplorer [3], and the MoDSE [42].
+Additionally, we implement ArchExplorer [2] for comparison with our ABA algorithm.
 The baselines represent recent SOTA work.
 All experiments are conducted on the Out-of-Order (OoO) CPU implemented by Gem5.
-The design space of the OoO processor is listed in Table 1.
+The design space of the OoO processor is listed in Table I.
 The values in the third column are formatted as “start number: end number: stride”.
@@
 The size of the design space exceeds $6.89\times 10^{35}$, making exhaustive search impractical.
-Table 1. Microarchitecture design space specification.
+TABLE I: Micro-architecture design space specification.
 Parameters Description
@@
 in $\mu$-ops
-8:48:4
+$8:4$
 Branch Predictor predictor type
@@
 The accuracy is mainly measured by Mean Absolute Percentage Error (MAPE), which calculates the average absolute percentage difference between real and predicted values. The formula for MAPE is:
-[…]
+ $\text{ MAPE }=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{\left|y_{i}-y_{i}^{\text{real }}\right|}{y_{i}^{\text{real }}}\right)\times 100.$ (5)
 where $n$ is the number of predicted design points, $y_{i}$ is the prediction objective value, i.e., IPC, power, or area in corresponding models, and $y_{i}^{real}$ is the real value; MAPE expresses the prediction error as a percentage of the actual values.
@@
 The PHV metric assesses the quality of the Pareto optimal set by quantifying the volume of the objective space dominated by non-dominated solutions. The PHV can be formulated as:
-(6) $\mathrm{PHV}(\Omega)=\int_{\mathrm{X}}\mathbb{1}[\boldsymbol{x}\prec\boldsymbol{x_{ref}}][1-\prod_{\boldsymbol{x}^{*}\in\Omega}\mathbb{1}[\boldsymbol{x}^{*}\nprec\boldsymbol{x}]]\mathrm{d}\boldsymbol{x}.$
+$\mathrm{PHV}(\Omega)=\int_{\mathrm{X}}\mathbb{1}[\boldsymbol{x}\prec\boldsymbol{x_{ref}}][1-\prod_{\boldsymbol{x}^{*}\in\Omega}\mathbb{1}[\boldsymbol{x}^{*}\nprec\boldsymbol{x}]]\mathrm{d}\boldsymbol{x}.$ (6)
 Where $\mathbb{1}(\textperiodcentered)$ is the indicator function, which outputs one if its argument is true and zero otherwise, specifically, if $x$ is the newly discovered design that is not dominated by other design points, then the PHV increases.
-An example is illustrated in Fig.2.
-7. Experiment Results
-7.1. Overall Analysis
+See Fig.
+2 for an example.
+VII Experiment Results
+VII-A Overall Analysis
 We evaluate the AttentionDSE framework using SPEC CPU 2017 workloads.
-7.1.1. Comparison on Pareto Optimal Set Exploration
-Figure 8. Comparison of PHV from IPC-Power optimization with SOTA DSE frameworks.
-Figure 9. Comparisons between the Pareto designs in (a) $Perf^{2}/(Power\times Area)$, (b) performance, (c) area, and (d) power.
+VII-A1 Comparison on Pareto Optimal Set Exploration
+Figure 8: Comparison of PHV from IPC-Power optimization with SOTAs.
+Figure 9: Comparisons between the Pareto designs in (a) $Perf^{2}/(Power\times Area)$, (b) performance, (c) area, and (d) power.
 Fig.
@@
 In this experiment, we use a random sampling method for design point search in MoDSE.
-Table 2. Detailed information of Pareto optimal set search.
+TABLE II: Detailed information of Pareto optimal set search.
 Methods
@@
 (Minutes)&(Ratio)
-BOOMExplorer 36.94 64 47 (9.4$\times$)
-MoDSE 37.04 87 16 (3.2$\times$)
-ArchExplorer 38.57 \ 828 (165.6$\times$)
-AttentionDSE 39.41 15 5 (1$\times$)
+[…]
+[…]
+[…]
+[…]
 Note: The experimental results are obtained from IPC-Power optimization on the 600.perlbench_s dataset.
-Table 2 shows the exploration results on the 600.perlbench_s dataset across the SOTA DSE frameworks.
-The “Time” represents the total exploration time of the DSE task.
-All DSE frameworks, except ArchExplorer, perform a hundred iterations for the Pareto optimal set search.
-The “Up to 99% PHV Iteration” metric indicates the convergence time of the search, where lower values are better.
-Experimental results demonstrate that AttentionDSE is 166 $\times$ faster than ArchExplorer, owing to ArchExplorer’s time-consuming instruction analysis method.
-Compared to regression-based models, AttentionDSE completes the exploration roughly 10$\times$ faster than BOOMExplorer.
-Even with MoDSE’s use of random sampling, AttentionDSE achieves exploration speeds approximately 3$\times$ faster.
-Additionally, AttentionDSE converges in fewer iterations, reducing iterations by 75% compared to BOOMExplorer and 83% compared to MoDSE, highlighting the efficiency of the ABA algorithm.
+[…]
 Overall, the results demonstrate that AttentionDSE provides significant improvements over existing work, validating the effectiveness of its novel methodologies in DSE.
-Figure 10.
-Comparison of MAPE with SOTA DSE frameworks.
+Figure 10: Comparison of MAPE with SOTA DSE frameworks.
 (a) MAPE of IPC prediction; (b) MAPE of power prediction; (c) MAPE of area prediction; (d) MAPE and normalized inference time with the increased number of parameters on the 600.perlbench_s dataset; (e) The training time changes with the increase in training set size on the 600.perlbench_s dataset.
-7.1.2. Comparison on Best Balanced Designs
+VII-A2 Comparison on Best Balanced Designs
 To study how high-performance designs balance PPA effectively, we select the designs with the highest performance for comparison.
@@
 9(d), leading to overall optimality across the design space.
-7.1.3. Comparison on Single-objective Predictor
+VII-A3 Comparison on Single-objective Predictor
 An accurate predictor is fundamental to an effective DSE framework.
@@
 10(a), (b), and (c) present the MAPE of AttentionDSE compared with three DSE frameworks.
+All the predictors are trained with one thousand design points.
 For example, AttentionDSE outperforms all workloads, achieving 37.1%, 33.4%, and 31.6% on average for IPC, Power, and Area prediction on MAPE, respectively, compared to the SOTA DSE framework (MoDSE).
 These results highlight the superiority of AttentionDSE over statistical regression methods.
-We also conduct experiments using an MLP regression model, as shown in Table 3, to demonstrate the superiority of the attention mechanism.
-Table 3. Comparison of AttentionDSE with the MLP model.
+We also conduct experiments using an MLP regression model, as shown in Table III, to demonstrate the superiority of the attention mechanism.
+TABLE III: Comparison of AttentionDSE with the MLP model.
+Methods
 […]
+MLP AttentionDSE
 IPC Power Area IPC Power Area
@@
 For AttentionDSE, the training time remains stable due to several aspects: first, the predictor in AttentionDSE consists of the minimal necessary components for an attention-based model, ensuring lower training time through its lightweight design; second, AttentionDSE adopts the PDA mechanism, further decreasing computational complexity.
-7.2. Optimization Effect Analysis
+VII-B Optimization Effect Analysis
 We further conduct an ablation experiment to demonstrate the effectiveness of the optimizations.
 We select 600.perlbench_s as the representative dataset, and the results are consistent across other datasets.
-7.2.1. Efficiency of Perception-driven Attention Mechanism
+VII-B1 Efficiency of Perception-driven Attention Mechanism
 To demonstrate the benefits of the PDA mechanism, we conduct experiments to evaluate the training time and memory footprint with varying numbers of parameters, highlighting the PDA mechanism’s superiority in high-dimensional DSE tasks.
 We set the model depth to ten and observe the impact on both training time and memory usage.
-Figure 11. The efficiency of PDA mechanism.
-Fig.
-11 shows the experiment results, the line chart illustrates the normalized training time, while the bar chart depicts the memory footprint.
-The training time for models without the PDA mechanism increases exponentially with the size of the design space.
-For example, when the design space expands sixfold, the training time increases by 30$\times$.
-In contrast, AttentionDSE with the PDA mechanism exhibits relatively stable memory usage as the number of parameters grows.
-This stability extends to large design spaces because the PDA mechanism only calculates within clusters that contain the most relevant parameters.
-Regarding memory footprint, in smaller design spaces, memory usage is primarily influenced by fixed data occupation during training.
-However, attention weights and other learnable parameters in larger design spaces gradually dominate memory usage.
-Figure 12. The impact of the PDA mechanism.
+Figure 11: The efficiency of PDA mechanism.
+[…]
+Figure 12: The impact of the PDA mechanism.
 Fig.
@@
 This improvement is primarily due to the enhanced understanding of local information within the pipeline stage provided by PDA.
-By focusing on parameter interactions within a defined window, PDA captures intricate relationships between parameters, resulting in more accurate predictions.
-7.2.2. Analysis on Attention-aware Bottleneck Analysis Algorithm
+By focusing on parameter interactions within a defined window, PDA captures relationships between parameters, improving prediction accuracy.
+VII-B2 Analysis on Attention-aware Bottleneck Analysis Algorithm
 The superiority of AttentionDSE derives from both its high-accuracy prediction model and the ABA algorithm.
 To more effectively illustrate the impact of the ABA algorithm and avoid interference from the inaccurate prediction model, we conduct experiments on AttentionDSE by substituting the ABA algorithm with other design point search methods.
-Due to the large design space in our experiment, exhaustive search methods like MoDSE become impractical.
-Additionally, methods similar to BOOMExplorer can only be applied to the GPR model.
-Therefore, we select the random sample search method as the baseline for comparison with the ABA algorithm.
+Given the large design space, exhaustive methods like MoDSE are impractical.
+Since BOOMExplorer-like methods apply only to GPR models, we use random sample search as the baseline to compare with ABA.
 Fig.
-13 visualizes the PHV curves and ADRS curves across iterations, comparing them with the random sample search method.
-Figure 13. The convergence of ABA algorithm for IPC-Power optimization.
-Regarding PHV convergence, the ABA algorithm notably achieves a higher PHV early on, consistently surpassing the random sample search method throughout various simulation iterations.
-Specifically, the ABA algorithm reaches a PHV of 38.2 within one hundred iterations, while the random sample search method requires more than 5$\times$ as many iterations to achieve similar levels.
-This efficiency results from the ABA algorithm’s effective analysis of parameter interactions and their impact on performance metrics, significantly reducing exploration time by 80% and guiding optimal parameter updates for enhanced performance.
+13 shows PHV and ADRS curves over iterations against this baseline.
+Figure 13: The convergence of ABA algorithm for IPC-Power optimization.
+[…]
 Concerning ADRS convergence, it serves to evaluate the quality of estimated Pareto optimal sets in multi-objective optimization by comparing them against the true Pareto set, reflecting the distribution of searched Pareto optimal design points.
@@
 In contrast, the random sample search method faces challenges in promptly reaching optimal intervals due to insufficient information on defects in current design points and challenges in selecting subsequent design points based on data distribution.
-8. Conclusion
-In this paper, we introduced AttentionDSE, a comprehensive DSE framework tailored for high-dimensional design spaces.
-AttentionDSE seamlessly synergizes the surrogate model with the acquisition function by incorporating the attention mechanism.
-The innovative PDA mechanism enhances prediction efficiency, while the ABA algorithm offers detailed insights into individual micro-architectural parameter contributions.
-These advancements improve the design process’s interpretability and minimize the need for expert knowledge.
-AttentionDSE represents the first attempt to apply attention mechanisms to DSE tasks, showcasing the potential of this approach in optimizing complex design spaces.
+VIII Conclusion
+In this paper, we propose AttentionDSE, the first attention-driven framework that unifies performance prediction and architectural exploration for high-dimensional CPU design.
+By leveraging transformer-based attention, it achieves accurate PPA estimation and exposes parameter-wise bottlenecks to guide efficient design refinement.
+Through PDA and ABA, AttentionDSE delivers faster convergence and better Pareto quality than prior methods.
+This work demonstrates the potential of attention mechanisms as a scalable and interpretable solution for automated micro-architecture design.